### v1

Abstract
Large Language Models (LLMs) have made notable progress in mathematical reasoning, yet they often rely on single-paradigm reasoning that limits their effectiveness across diverse tasks. In this paper, we introduce Chain-of-Reasoning (CoR), a novel unified framework that integrates multiple reasoning paradigms—Natural Language Reasoning (NLR), Algorithmic Reasoning (AR), and Symbolic Reasoning (SR)—to enable synergistic collaboration. CoR generates multiple potential answers using different reasoning paradigms and synthesizes them into a coherent final solution. We propose a Progressive Paradigm Training (PPT) strategy that allows models to progressively master these paradigms, culminating in the development of CoR-Math-7B. Experimental results demonstrate that CoR-Math-7B significantly outperforms current SOTA models, achieving up to a $41.0\%$ absolute improvement over GPT-4 in theorem proving tasks and a $7.9\%$ improvement over RL-based methods in arithmetic tasks. These results showcase the enhanced mathematical comprehensive ability of our model, achieving significant performance gains on specific tasks and enabling zero-shot generalization across tasks.
Figure 1: A comprehensive comparative analysis of CoR-Math-7B and baseline models across challenging mathematical tasks. (a) demonstrates the effectiveness of CoR-Math-7B (zero-shot) in theorem proving tasks, specifically emphasizing its advantage over baseline and their fine-tuned variants over SFT and RL. (b) shows a resource-efficiency analysis for arithmetic computation tasks, plotting model performance vs. SFT data size, in which CoR-Math-7B attains the optimal resource efficiency, achieving near-optimal zero-shot performance.
1 Introduction
While LLMs have demonstrated strong performance in solving mathematical tasks (Feigenbaum et al., 1963; Hosseini et al., 2014), advanced open-source reasoners still struggle with solving comprehensive mathematical problems, including both arithmetic computation and theorem proving.
Figure 2: The reasoning process under different paradigms: (a) In single-paradigm reasoning, each reasoning step relies on the same knowledge medium, such as Natural Language (NL), algorithms, or symbols. (b) In tool-integrated single-paradigm, NL is used for reasoning, while code assists in solving specific sub-problems. After obtaining the execution results, the reasoning continues using NL. (c) The proposed CoR reasoning framework, along with several examples, shows that multi-paradigm reasoning allows for varying reasoning depths to address different types of problems.
Existing works (Xin et al., 2024; Yang et al., 2024; Wu et al., 2024) are often trained on specific tasks, aiming to enhance their ability to independently derive answer based on specific structured knowledge representation. This representation is known as the reasoning paradigm, involving Natural Language Reasoning (NLR), Algorithmic Reasoning (AR), and Symbolic Reasoning (SR), as illustrated in Figure 2 (a). Specifically, NLR leverages natural language text for reasoning based on human common sense and semantic context, with explicit step-by-step natural language explanations (Wei et al., 2022). AR leverages code to focus on the computer’s operations and execution process, performing inference on the final target, such as the generation of Python code for compilation (Chen et al., 2023; Gao et al., 2023) to obtain the final answer. SR utilizes logical symbols and axiomatic systems as the fundamental tools for rigorously formalized reasoning, with current methods (Xin et al., 2024; Huang et al., 2024; Wu et al., 2024) considering numerous symbolic trajectories through a tree-based search process for theorem proving. However, these methods focus on improving single-paradigm reasoning and overlook the potential of collaboration among multiple paradigms, which restricts their single-task performance and hinders their cross-task generalization.
Researchers have explored various strategies to tackle these challenges. To improve the single-task performance, some works have explored integrating tools to overcome the limitations of single-paradigm reasoning (Gou et al., 2024; LI et al., 2024), as shown in Figure 2 (b). These methods combine natural language with code-based algorithms (tools) to enable interleaved reasoning and generate a final answer. Although these methods acknowledge the potential benefits of integrating different reasoning paradigms, they often rely on a single paradigm to complete the reasoning process. They overlook the possibility that the second paradigm could independently complete the reasoning, thereby constraining its overall potential. Besides, in an effort to overcome the challenge of achieving cross-task generalization, several studies (Shao et al., 2024; Huang et al., 2024) incorporate data samples from diverse task types into large-scale training datasets, such as those drawn from theorem proving tasks that focus exclusively on SR solutions, or from arithmetic problems that emphasize NLR solutions. Although models trained on such data are capable of cross-task reasoning, they still rely on demonstrations for effective transfer.
To address these limitations, we propose a novel unified reasoning framework, Chain-of-Reasoning (CoR), which chains NLR, AR, and SR together to generate synergistic benefits. As illustrated in Figure 2 (c), CoR enables multi-paradigm reasoning for a given problem, which applies different reasoning paradigms to derive multiple potential answers and are then summarized into a final answer. The framework allows the model to continue reasoning using additional paradigms based on previously generated ones, facilitating collaboration among paradigms to enhance individual task performance. Moreover, CoR implements unified multi-paradigm reasoning across different tasks to obtain the required answers, thereby achieving zero-shot reasoning across tasks. In detail, by adjusting the prompts, the depth of reasoning can be varied, which improves the model’s adaptability to diverse tasks. By constructing such multi-paradigm reasoning paths, we create a training dataset named MPM, comprising $167$k reasoning paths. Furthermore, we propose a novel Progressive Paradigm Training (PPT) approach, which enables models to progressively master multiple reasoning paradigms, resulting in CoR-Math-7B.
We evaluated CoR-Math-7B on three representative and challenging mathematical reasoning benchmarks that effectively demonstrate comprehensive reasoning capabilities, encompassing both arithmetic computation and theorem proving. These benchmarks include MATH (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), and MiniF2F (Zheng et al., 2022). Our results show that LLMs equipped with CoR significantly outperform current SOTA baselines. In the theorem proving task as shown in Figure 1 (a), CoR-Math-7B surpasses GPT-4o’s (Hurst et al., 2024) zero-shot performance by a remarkable absolute $41.0\%$ and outperforms all other few-shot reasoners. In arithmetic tasks, our CoR-Math-7B first performs NLR and SR as a rehearsal, which enhances the accuracy of AR and demonstrates the effectiveness of synergy. For example, CoR-Math-7B achieves an absolute improvement of $7.9\%$ over Reinforcement Learning (RL) methods (Shao et al., 2024) that utilize the same base model on the MATH dataset. Furthermore, as shown in Figure 1 (b), CoR-Math-7B efficiently utilizes resources to surpass the optimal performance curve of single-paradigm approaches. Compared to mainstream methods that perform extensive searches within a single paradigm, our CoR framework achieves better results by extending test-time inference across multiple paradigm spaces. These results show that our method achieves comprehensive mathematical problem solving abilities though multi-paradigm reasoning, requiring fewer training data and reducing reasoning costs.
2 Related Work
Reasoning Paradigms in LLMs. Recent advancements in LLMs have primarily concentrated on enhancing single-paradigm reasoning, where each paradigm serves as a distinct medium for knowledge representation and logical inference. NLR utilizes human-readable language to solve commonsense reasoning and step-by-step deductions (Wei et al., 2022; Yao et al., 2023; Zhou et al., 2023; Besta et al., 2024; Sel et al., 2024). For precise numerical computations, AR is to generate executable code, such as Python scripts, to perform accurate calculations through execution (Chen et al., 2023; Rozière et al., 2023). In the realm of theorem proving, SR leverages logical symbols and axiomatic systems to ensure rigorous and verifiable proofs (Xin et al., 2024). Furthermore, some approaches leverage external tools within the single-paradigm to enhance reasoning capabilities (Gou et al., 2024), such as using calculators or code interpreters, while maintaining the core reasoning process within a single paradigm. While these paradigms excel within their specific areas, they often struggle with cross-domain generalization and dynamic environments. To address these issues, the CoR framework facilitates collaboration between various reasoning paradigms to enable zero-shot multitasking generalization.
Mathematical Problem Solving with LLMs. Advanced research demonstrates the significant potential of LLMs in solving mathematical problems (Zhu et al., 2023). Several studies have developed unified mathematical solvers by synthesizing mathematical data, capable of addressing various challenges such as arithmetic calculations and theorem proving (Huang et al., 2024; Shao et al., 2024). However, recent research tends to focus on optimizing single paradigms. For instance, in the field of arithmetic computation, there is a tendency to incorporate programming tools to aid natural language mathematical reasoning (Gou et al., 2024). In theorem proving, researchers not only utilize specialized annotated data, including pre-training data and reinforcement learning reward data, but also employ tree-based search methods to generate numerous possible solutions (Ying et al., 2024b; Xin et al., 2024). These methods either fail to fully consider the importance of complete reasoning or rely heavily on large-scale search within the solution space to fully explore the performance improvement potential of a single paradigm. To address these issues, we introduce the MPM dataset and present the CoR-Math-7B model, trained on it, which leverages the synergistic effects of multi-paradigm reasoning to navigate towards the optimal solution within the expanded multi-paradigm solution space.
Figure 3: An overview of (a) the Multi-Paradigm Math (MPM) dataset construction process, involving reconstruction, extension, and theorem prover verification, and (b) the Progressive Paradigm Training (PPT) method, where the model is trained with increasing reasoning paradigms in stages.
3 Chain-of-Reasoning Framework
3.1 Overview
Our CoR framework aims to enable LLMs to perform a series of multi-paradigm reasoning on any type of mathematical problem, ultimately arriving at a solution. Specifically, given a mathematical problem $x$, LLMs ($\mathbb{P}$) can infer the result $y$ by following multiple reasoning paradigms, where each reasoning paradigm $\tau$ includes $n$ reasoning paths $\{rp_{1},...,rp_{n}\}$. To formally represent this single-paradigm scenario, we can formulate it as $y\sim\mathbb{P}(y|x,\tau)$. To simplify the process for each reasoning paradigm, we set $n=1$ as default (Details in Section 6).
Inspired by recent works (Wei et al., 2022) in encouraging step-by-step reasoning, we introduce CoR, which extends from a single paradigm to three paradigms $\Gamma=(\tau_{1},\tau_{2},\tau_{3})$. For generating multiple chained reasoning paradigms for a given problem, CoR follows the steps outlined below. The reasoning process begins with the initialization of the problem $x$ and the first reasoning paradigm $\tau_{1}$. Subsequently, each paradigm $\tau_{i}$ in the sequence $\Gamma$ is generated based on the problem $x$ and the previously established paradigms $\tau_{1},...,\tau_{i-1}$, as represented by $\tau_{i}\sim\mathbb{P}(\tau_{i}|x,\tau_{1},...,\tau_{i-1})$. Finally, the outcomes from all reasoning paradigms are aggregated to derive the final result $y$, expressed as $y\sim\mathbb{P}(y|x,\tau_{\text{NLR}},\tau_{\text{SR}},\tau_{\text{AR}})$, considering three paradigms: NLR, SR, and AR. In detail, $\tau_{\text{SR}}$ and $\tau_{\text{AR}}$ include both the complete reasoning paths and the interaction processes with tool outputs. In our experiments, $\tau_{\text{SR}}$ utilizes the Lean prover, and $\tau_{\text{AR}}$ leverages a Python compiler as the algorithmic interactive tool. CoR possesses the capacity for broad application in complex reasoning scenarios. In the context of mathematical reasoning, it manifests as CoR-Math, with potential for future adaptations such as CoR-Code.
According to this framework, as shown in Figure 3, the training pipeline consists of three key phases: (a) collecting a Multi-Paradigm Mathematical (MPM) dataset, with deep reasoning paths based on multiple reasoning paradigms; (b) introducing a Progressive Paradigm Training (PPT) strategy to shape the model’s reasoning behavior, gradually mastering more reasoning paradigms; and (c) leveraging the trained LLM for zero-shot inference, allows multi-paradigm reasoning with adaptable depth, and uses Sequential Multi-Paradigm Sampling (SMPS) to explore a diverse solution space.
3.2 Collecting Dataset
[…]
Stage 1: Reconstructing and Extending. We introduce a universal text template designed for multi-paradigm reasoning (Details in Section A.1). This template standardizes the positioning of distinct reasoning paradigms and defines the relationships between them. It accommodates various reasoning depths and supports flexible combinations of different reasoning paradigms. As shown in Figure 2 (c), Problem 1 demonstrates an instance incorporating both NLR and SR paradigms, while Problem 2 showcases an instance that integrates three reasoning paradigms. To ensure data integrity and prevent potential biases, we pre-process the Numina-TIR (LI et al., 2024) and Lean-Workbook (Ying et al., 2024a) datasets through a two-step process. First, samples without corresponding solutions are removed. Second, we further reconstruct and extend these datasets by leveraging powerful LLMs $\mathcal{G}$, such as GPT-4o (Hurst et al., 2024). These models generate missing reasoning paradigms $\tau_{g}$ and refine existing ones $\tau^{\prime}$, ensuring comprehensive coverage and logical consistency. To effectively guide the processes of augmentation and refinement, we develop tailored prompts $p_{s}$ for each seed dataset. Examples of these prompts are detailed in Section A.2. The procedure can be described in:
$\tau_{g}\sim\mathbb{P}_{\mathcal{G}}(\tau_{g}\mid p_{s}\oplus x\oplus y\oplus\tau^{\prime}),\tau^{\prime}\in\{\tau_{\text{NLR}},\tau_{\text{SR}},\tau_{\text{AR}}\},$ (1)
where $\oplus$ means concatenation. After that, we conduct a manual review of all samples. Given that alternative approaches are readily verified through external tools, we focus on the accuracy of the NLR and AR. This approach considerably lowers the requisite skill level of annotators. Furthermore, to avoid data leakage, we calculate the Levenshtein distance (Miller et al., 2009) between problems in the training and test samples, while details are provided in the Section B.1. As a result, this phase yields the MPM-raw dataset, comprising approximately $285,000$ synthetic samples.
Stage 2: Revising. The MPM-raw dataset interacts with the Lean prover to verify proof steps, guiding the filtering and modification of reasoning paths. The process submits the proof $\tau_{\text{SR}}$ to the prover. Once the prover successfully completes the proof without returning any errors, the entire multi-paradigm reasoning path is collected to the MPM dataset. Otherwise, the error information $\varepsilon$ returned by the prover is fed into a revising model $\mathbb{P}_{R}$. Furthermore, this model generates a revised proof $\tilde{\tau}_{\text{SR}}$ based on a prompt $p_{\varepsilon}$. The relationship can be expressed as:
$\tilde{\tau}_{\text{SR}}\sim\mathbb{P}_{R}(\tilde{\tau}_{\text{SR}}\mid p_{\varepsilon}\oplus x\oplus y\oplus\tau_{\text{SR}}),$ (2)
where the revised proof $\tilde{\tau}_{\text{SR}}$ is then resubmitted to the Lean prover for verification. This iterative process continues for a maximum of $64$ iterations or until the prover verifies $\tilde{\tau}_{\text{SR}}$ as correct. In detail, we utilize DeepSeek-Prover-V1.5 (Xin et al., 2024) as the revising model.
As a result, the MPM dataset comprises $82,770$ problems and $167,412$ multi-paradigm reasoning solutions.
3.3 Training
Inspired by recent advances (LI et al., 2024), we introduce the Progressive Paradigm Training (PPT) strategy, which enables LLMs to gradually master an increasing variety of reasoning paradigms. As shown in Figure 3 (b), this training framework enables the model to progressively expand its reasoning capability by introducing different types of reasoning data in stages. Each training stage uses a different combination of reasoning paradigms. In the stage ➀, given the dominance of NL in the pre-training data of language models (Dubey et al., 2024), we create Numina-CoT as an initialized teaching stage. This is done by modifying the original Numina-CoT dataset (LI et al., 2024) according to our universal text template, which enables the model to learn to use NL to solve complex mathematical problems. Based on the question $x$, the model performs reasoning $\tau_{\text{NLR}}$ and generates the answer $y$. The generated sequence is $z=[x]\tau_{\text{NLR}}y$, where $[x]$ represents the tokens of the inputs. For simplicity, we first consider the loss function for a single sample:
$\mathcal{L}_{\text{sample}}=-\sum_{t=1}^{|z|}\log\mathbb{P}_{\theta}(z_{t}\mid z_{<t}),$ (3)
where $\theta$ represents the model parameters, $z_{t}$ is the $t$-th token in the generated sequence, and $z_{<t}$ indicates all tokens before the $t$-th token in the generated sequence. In stage ➁, considering a certain proportion of code corpora in the pre-training data, we expand the training dataset to include two paradigms: NLR and AR. Similar to Numina-CoT, we modify the original Numina-TIR to create Numina-TIR. With this modified dataset, the generated sequence is $z=[x]\tau_{\text{NLR}}\tau_{\text{AR}}y$. After this stage, the model can handle problems that require precise answers. Finally, in the stage ➂, we further expand the training data to three reasoning paradigms by utilizing the MPM dataset, where $z$ is $[x]\tau_{\text{NLR}}\tau_{\text{AR}}\tau_{\text{SR}}y$. After full PPT stages, the trained CoR-Math-7B model not only masters NLR and AR but also can perform rigorous logical SR.
3.4 Inference
We introduce an inference method that combines variable reasoning depth and sequential multi-paradigm sampling to solve comprehensive mathematical tasks, enabling the model to adapt its reasoning process based on task-specific requirements.
Prompts with variable reasoning depth. In the zero-shot inference phase, the CoR-Math-7B exhibits proficiency in multi-paradigm reasoning. This proficiency allows for the implementation of multi-paradigm reasoning with adaptable reasoning depths, achieved by adjusting prompts according to the specific demands of a given task. Initially, the model is prompted to conduct NLR, thereby activating a wide range of knowledge patterns associated with natural language from the pre-training corpus. Subsequently, the inference method is tailored to the specific problem type. For instance, in the context of theorem proving, the model transitions to SR, which is better suited for formal deduction. Given that the output of the model during SR is structured in Lean 4, the relevant proof segment can be extracted as the final solution. In cases of arithmetic computation, the model first performs NLR, followed by SR to strengthen logical coherence, and ultimately employs AR for precise calculations. A summary box is utilized to present the final results. This adaptable paradigm effectively captures the distinct reasoning patterns inherent in different paradigms and demonstrates flexibility in accommodating a variety of scenarios.
Sequential Multi-Paradigm Sampling (SMPS). Instead of token-level sampling based on tree structures (Qiu et al., 2024; Xiong et al., 2024) within single-paradigm reasoning paths, we purpose sequential paradigm-level sampling method, named SMPS. This approach allows the model to generate outputs by sequentially sampling across different paradigms. For instance, in a two-paradigm reasoning scenario, the model first instantiates $J$ distinct paths for the initial reasoning paradigm $\tau_{1}$.
$\tau_{1j}\sim\mathbb{P}(\tau_{1j}\mid x),\quad\forall j\in{1,\dots,J}$ (4)
Subsequently, for each of these $\tau_{1}$ paths, the model instantiates $K$ possible paths for the secondary reasoning paradigm $\tau_{2}$
$\tau_{2k}\sim\mathbb{P}(\tau_{2k}\mid x,\tau_{1j}),\quad\forall k\in{1,\dots,K},\forall j$ (5)
This hierarchical sampling process yields a total of $J\times K$ potential responses, denoted as $y$.
$y_{jk}\sim\mathbb{P}(y_{jk}\mid x,\tau_{1j},\tau_{2k}),\quad\forall j,k$ (6)
In summary, the SMPS method leverages the combinatorial expansion of reasoning paths to explore a diverse paradigm-based solution space comprehensively. This structured approach enhances the robustness and depth of the reasoning results.
4 Experimental Settings
4.1 Evaluation Setup
Datasets. To evaluate our model’s mathematical reasoning capabilities comprehensively, we utilize a suite of widely recognized and representative benchmark datasets. These datasets encompass a diverse range of mathematical reasoning tasks, including arithmetic computation and theorem proving. The arithmetic computation ability is evaluated on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) datasets. These datasets span a spectrum of difficulty levels, from elementary arithmetic to mathematical competitions. Theorem proving ability is evaluated on the MiniF2F test set (Zheng et al., 2022), which features mathematical problems of Olympiad difficulty.
Metrics. Accuracy serves as the primary metric for performance evaluation across all datasets. In detail, we follow the widely-used CoT settings (Wei et al., 2022) to obtain solutions for arithmetic problem tasks, with numerical answers rounded to the nearest integer. The SymPy library [^1] https://www.sympy.org/ is utilized for parsing and evaluating mathematical expressions. To address potential variations in numerical representation, the model is instructed to explicitly state the final answer in the conclusion, following the approach in Numina models (LI et al., 2024). For theorem proving datasets, we follow the recent advances (Xin et al., 2024) to adapt the publicly available MiniF2F benchmark for the Lean 3 theorem proving environment to the Lean 4 environment [^2] https://github.com/rahul13613/miniF2F-lean4. These adjustments ensure the use of the most up-to-date formal verification tools. The pass@$N$ metric evaluates the capacity of the model to generate correct proofs within $N$ attempts. Specifically, our CoR-Math-7B employs SMPS, which operates by performing NLR followed by SR. Therefore, the sample budget $N$ is $N_{\text{NLR}}\times N_{\text{SR}}$. Detailed settings of all models are provided in Section A.5.
4.2 Implementation Details
We fine-tuned widely-used DeepSeek-Math-Base 7B (Shao et al., 2024) and Llama-3.1 8B (Dubey et al., 2024) models, employing our PPT method on MPM dataset. Unless otherwise specified, CoR-Math-7B model is based on DeepSeek-Math-Base 7B. The details are available in Section A.3.
4.3 Baselines
This study examines three categories of baseline models, with results reported using CoT prompting.
General-purpose mathematical models: Considering the general applicability of our CoR-Math-7B model, we present the SOTA general-purpose mathematical models, including Mustard (Huang et al., 2024), DeepSeek-Math (Shao et al., 2024), InternLM-Math (Ying et al., 2024b), Llama-3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Llemma (Azerbayev et al., 2024).
Task-specific mathematical models: We consider several expert models on mathematical optimization for specific tasks. For expert models on arithmetic computation, we include several models that undergo specialized reinforcement in arithmetic computation to achieve enhanced performance, such as more data and tool invocation. These models encompass Qwen2.5-Math (Yang et al., 2024), WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2024), DART-Math (Tong et al., 2024), InternLM-Math (Ying et al., 2024b), , DeepSeek-Math-Instruct / RL (Shao et al., 2024), Xwin-Math (Li et al., 2024), ToRA (Gou et al., 2024), and NuminaMath (LI et al., 2024). For expert models on theorem proving, we consider several SOTA models with substantial improvements, including reinforcement learning and Monte Carlo search. We include LLM-Step (Welleck and Saha, 2023), GPT-f (Polu and Sutskever, 2020), Lean-STaR (Lin et al., 2024), Hypertree Proof Search (Lample et al., 2022), DeepSeek-Prover (Xin et al., 2024), and InternLM2.5-StepProver (Wu et al., 2024).
Foundation models and proprietary models: We introduce current advanced open-source foundational models, which are Llama-3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Mixtral (Jiang et al., 2024). We also introduce proprietary models, which are GPT-4 (OpenAI, 2023), GPT-4o (Hurst et al., 2024) and o1-mini (OpenAI, 2024), as robust competitors.
5 Results and Analysis
5.1 Main Results
This section presents a comprehensive evaluation of CoR-Math-7B’s capabilities in diverse mathematical reasoning tasks. We systematically assess its performance against SOTA models and expert models on specific domains using a zero-shot evaluation setting. Our objective is to demonstrate how multi-paradigm approach enables CoR-Math-7B to achieve superior general mathematical reasoning abilities and computational efficiency, compared to methods based on a single reasoning paradigm.
5.1.1 Comparisons with General-purpose Mathematical Models
Table 1: A overall comparison of CoR-Math-7B with three types of general mathematical reasoners (proprietary, foundational, and general-purpose mathematical models (GMM)) on three mathematical benchmarks. Results are shown for zero-shot (denoted by $\emptyset$) or few-shot settings by default. For the MiniF2F benchmark, we report the best results from the relevant literature with a specified sample budget. Bolded scores are the best performance among all models except for the proprietary one.
Model Model Types MATH GSM8k miniF2F-test Sample Budget ($N$) miniF2F-test Pass@1 Pass@1 Pass@$N$ o1-mini Proprietary 90.0 94.8 1 13.2 GPT-4 Proprietary 42.5 87.1 128 24.6 GPT-4o Proprietary 76.6 90.5 128 25.0 Llama-3.1-8B Foundation 4.2 6.2 128 25.8 Llama-3.1-8B-Instruct Foundation 47.2 76.6 128 23.4 Mistral-7B Foundation 14.3 40.3 1$\times$32$\times$100 22.1 Mixtral-8x7B Foundation 28.4 74.4 1$\times$32$\times$100 23.4 MUSTARD GMM 13.8 27.9 1 7.8 Llemma-7B GMM 18.6 41.0 1$\times$32$\times$100 26.2 DeepSeek-Math-7B-Base GMM 11.8 22.2 1$\times$32$\times$100 28.3 InternLM2-Math-7B-Base GMM 21.5 49.2 1$\times$32$\times$100 30.3 InternLM2-Math-Plus-7B GMM 53.0 85.8 1$\times$32$\times$100 43.3 CoR-Math-7B GMM 66.7 88.7 128$\times$1 52.9 32$\times$100 59.4 128$\times$128 66.0
To explore the versatility of CoR-Math-7B and assess its comprehensive mathematical reasoning capabilities, we compare it with widely-used models and SOTA general mathematical models. As shown in Table 1, CoR-Math-7B achieves the best performance across three challenging benchmarks in a zero-shot setting. These results collectively highlight the strong comprehensive mathematical reasoning abilities of our model. The findings are detailed below: (1) arithmetic computation task: our CoR-Math-7B achieves optimal results across all arithmetic computation benchmarks. For example, it exceeds InterLM2-Math-Plus-7B by a $13.7\%$ absolute improvement on the MATH dataset. (2) theorem proving task: our CoR-Math-7B achieves the best zero-shot result on MiniF2F, even surpassing the proprietary GPT-4o model by a $41\%$ absolute increase in few-shot setting. (3) zero-shot vs few-shot: The zero-shot results of CoR-Math-7B outperform all few-shot results from other models. For instance, the zero-shot performance of CoR-Math-7B demonstrates a $14.0\%$ absolute improvement over the few-shot performance of DeepSeek-math-7B-Base on the MiniF2F dataset. When testing on the MiniF2f dataset revealed limited proficiency in Qwen2.5-Math (Yang et al., 2024) in Lean 4, indicating a focus on arithmetic computation over theorem proving. Therefore, we classify Qwen2.5-Math as an arithmetic computation model rather than a general mathematical model and compare it in Section 5.1.3. These findings indicate the strong generalization capability and comprehensive mathematical reasoning ability of the CoR framework. This suggests NLR descriptions and SR verification rehearse precise mathematical reasoning for AR. In summary, the results indicate that the CoR framework can handle a diverse range of complex reasoning challenges, and leveraging the synergy between different paradigms can enhance general reasoning capabilities.
5.1.2 Comparisons with Expert Models on Theorem Proving
Table 2 shows CoR-Math-7B’s exceptional zero-shot performance on the MiniF2F, achieving $66.0\%$ accuracy without demonstrations while competing methods require few-shot examples. Its zero-shot performance is particularly noteworthy when considering computational efficiency.
Table 2: The zero-shot ($\emptyset$) performance of CoR-Math-7B and the few-shot performance of optimized models on the MiniF2F benchmark. These models incorporate optimizations related to theorem proving. The highest scores are in bold.
Model Sample Budget $N$ MiniF2F LLMStep 1$\times$32$\times$100 27.9 GPT-f 64$\times$8$\times$512 36.6 Hypertree Proof Search 64$\times$5000 41.0 Lean-STaR 64$\times$1$\times$50 46.3 DeepSeek-Prover-V1.5-Base 6400 42.2 DeepSeek-Prover-V1.5-SFT + RMaxTS 4$\times$6400 56.3 DeepSeek-Prover-V1.5-SFT + RMaxTS 32$\times$6400 60.2 DeepSeek-Prover-V1.5-RL + RMaxTS 4$\times$6400 59.6 DeepSeek-Prover-V1.5-RL + RMaxTS 32$\times$6400 63.5 InternLM2.5-StepProver 64$\times$32$\times$100 54.5 InternLM2.5-StepProver-BF 1$\times$32$\times$600 47.3 InternLM2.5-StepProver-BF 256$\times$32$\times$600 59.4 InternLM2.5-StepProver-BF+CG 2$\times$32$\times$600 50.7 InternLM2.5-StepProver-BF+CG 256$\times$32$\times$600 65.9 CoR-Math-7B 128$\times$128 66.0
For instance, under similar computational constraints, our CoR-Math-7B surpasses DeepSeek-Prover-V1.5-RL + RMaxTS by $6.4\%$ absolute points ($66.0\%$ vs $59.6\%$ with $128\times 128$ vs $4\times 6400$) and InternLM2.5-StepProver-BF+CG by $15.3\%$ absolute points ($66.0\%$ vs $50.7\%$ with $128\times 128$ vs $2\times 32\times 600$). Compared to significantly larger sample sizes, CoR-Math-7B retains its competitive edge. For example, zero-shot CoR-Math-7B outperforms few-shot InternLM2.5-StepProver-BF+CG by $0.1\%$ in absolute score, despite considering approximately $300$ times more potential solutions. These results highlight the effectiveness of multi-paradigm reasoning and demonstrate how CoR efficiently searches paradigm-based solution spaces.
5.1.3 Comparisons with Expert Models on Arithmetic Computation
Table 3 shows the significant competitiveness of our CoR-Math-7B when compared to expert models on arithmetic computation. Compared to methods like ToRA and NuminaMath that use code as a tool without full reasoning, our CoR-Math-7B with multi-paradigm reasoning outperforms them on all benchmarks. For instance, on the competition-level MATH dataset, CoR-Math-7B achieves an $11.4\%$ absolute improvement over the tool-integrated NuminaMath-7B-TIR. This reinforces the effectiveness of complete code-based reasoning compared to the fragmented interleaving of natural language and Python code.
To further explore the relationship between resource utilization (training sample size) and performance, we fit a quadratic function to the current SOTA models and conduct a Pareto frontier analysis (Branke et al., 2008). Based on Table 3, we generate the fitting Pareto curve shown in Figure 1 (b), which is based on all reasoning methods. Our CoR-Math-7B achieves superior performance compared to the optimal performance of single-paradigm reasoning methods with similar data volume. This suggests the existence of a new optimal curve based on multi-paradigm reasoning methods, which outperforms the curve of single-paradigm approaches.
[…]
5.2 Ablation Study
We conduct a series of ablation studies, providing insights into the key factors that drive the performance of our multi-paradigm reasoning framework.
5.2.1 Impact of Stages in PPT Method
Figure 4: An evaluation of the effectiveness of the PPT strategy on mathematical reasoning tasks. We present the zero-shot Pass@1 results on the MATH and GSM8k benchmarks, across three cumulative stages of the PPT strategy based on the DeepSeek-Math and Llama-3.1 models. . The results highlight the cumulative effectiveness of the PPT strategy, showing increased performance for both models across all benchmark datasets with each progressive stage. These results indicate that the PPT strategy is a viable method to enhance mathematical reasoning capabilities.
To evaluate the effectiveness of PPT methods, we employ two base models, Llama-3.1-8B (Dubey et al., 2024) and Deepseek-Math-7B-Base (Shao et al., 2024). We present the results of the PPT method across different stages on these two models for the MATH and GSM8K benchmarks. As shown in Figure 4, the vanilla models exhibit limited performance on both benchmarks. After stage ①, the performance of all models improves significantly. For instance, Llama-3.1-Base achieves absolute gains of $47.9\%$ on MATH and $61.0\%$ on GSM8K. This demonstrates the necessity of understanding mathematical concepts and warming up the models. Subsequently, we conduct stage ② training, which utilizes data with NLR and AR. The limited performance gains are potential because the pre-training data already includes code and natural language corpus. Both stage ① and stage ② primarily reactivate the base models in the mathematical domain, and stage ① already incorporates NLR training. In the final stage ③, we consider training with three paradigms, and all models experience performance enhancements. This suggests that introducing rare or unseen paradigms helps to improve the reasoning ability of learned paradigms.
5.2.2 Impact of Reasoning Paradigms Order
We investigated the impact of varying the sequence of reasoning paradigms employed by CoR-Math-7B during zero-shot inference for arithmetic computation tasks. Our experiments involved modifying only the prompt structure to enforce different paradigm orders. Given that NLR closely aligns with the pre-training of LLMs, it was consistently positioned as the initial paradigm in our experiments. Table 4 presents the accuracy for the MATH and GSM8K benchmarks under two different paradigm orders. The results indicate that performing SR before AR achieves the best performance. Specifically, on the MATH benchmark, this order achieves an accuracy of $66.7\%$, an improvement over the $49.9\%$ attained by the sequence that AR before SR. A possible explanation for this observation lies in the synergistic effect of the ordered paradigms. SR plays a crucial role in decomposing the problem into manageable sub-steps, thereby providing a structured foundation for subsequent AR. By placing SR before AR, we hypothesize that the model is better equipped to perform accurate calculations, leveraging the intermediate reasoning steps generated by SR. Furthermore, positioning AR immediately preceding the Summary phase may facilitate a more coherent integration of the computational outcome into the final answer.
Table 4: Zero-shot Pass@1 results for varying paradigm orders on the MATH and GSM8k benchmarks. The best results are in bold.
[…]
MATH 66.7 49.9
GSM8K 88.7 84.2
5.2.3 Impact of Scaling Model Size
Table 5: Zero-shot Pass@1 results on the MATH and GSM8k benchmarks across different model scales. The best scores are in bold.
Model MATH GSM8k
Qwen2.5-Math-1.5B (CoR) 57.6 84.5
Qwen2.5-Math-7B (CoR) 64.7 90.0
Table 5 investigates the performance of various base models with different parameters under the CoR framework. We employ the Qwen2.5-Math models (Yang et al., 2024) with $1.5$ billion and $7$ billion parameters as the base model. Results demonstrate that our CoR framework achieves superior performance with increasing parameter scale. Specifically, when comparing a model with $7$ billion parameters to one with $1.5$ billion parameters, the former attained an absolute increase of $7.1\%$ on the MATH benchmark and $5.5\%$ on the GSM8K benchmark. This evidence supports the claim that CoR exhibits scalability with respect to model parameter scaling.
6 Discussion on Reasoning Hierarchy: Paradigms, Paths, and Steps
This section attempts to explore the essence of reasoning, contrasts CoR with current methods, examines the factors contributing to its effectiveness, and provides a new perspective for future research.
Figure 5: An overview of the reasoning hierarchical structure. (a) A single reasoning paradigm depicts multiple distinct reasoning paths. (b) An example of single-paradigm reasoning includes one reasoning path, which contains several reasoning steps. (c) Multi-paradigm reasoning includes several distinct reasoning paradigms.
As shown in Figure 5, this paper posits that reasoning texts generated by LLMs exhibit a reasoning hierarchical structure, which consists of three levels: reasoning paradigms, reasoning paths, and reasoning steps.
  • Reasoning steps represent the fundamental units, each comprising one or more tokens and encompassing an incomplete stage of the solution process.
  • Reasoning paths consist of several reasoning steps, forming a complete line of reasoning that typically includes a final answer and the solution process.
  • Reasoning paradigms comprise one or more reasoning paths. They often contain multiple potential final answers, thus necessitating a summarization method, such as a summary module, to extract the ultimate answer. Furthermore, a reasoning paradigm utilizes a single knowledge media, such as natural language.
Figure 6: A comparison of reasoning paradigms in advanced studies. (a) deep reasoning, characterized by serial concatenation of reasoning paths. (b) interleaved reasoning, which integrates secondary methods for guidance within a dominant paradigm. (c) multi-paradigm reasoning (Our proposed CoR), a framework designed to harness the synergistic effects between different reasoning paradigms.
Furthermore, current studies can be categorized based on their focus. As shown in Figure 5 (a), contemporary work concentrates within a single paradigm, optimizing along two dimensions: depth (the number of reasoning steps) and width (the number of reasoning paths). For instance, regarding reasoning depth, the CoT method (Wei et al., 2022) employs prompts to increase intermediate steps within one reasoning path to achieve higher performance. Concerning reasoning width, some approaches (Wang et al., 2023) involve altering sampling techniques to generate multiple distinct reasoning paths. Random sampling exemplifies this. Other studies (Zhu et al., 2023) employ scoring mechanisms for reasoning steps, guiding LLMs to generate several more desirable reasoning paths. Monte Carlo search illustrates this. Building upon these generated reasoning paths, existing work proposes various integration strategies to obtain the final answer, such as Best-of-N and Self-consistency (Wang et al., 2023). Specifically, as shown in Figure 6 (a), recent advanced studies introduce the deep reasoning method, which focuses on generating serial concatenation of reasoning paths followed by summarization (like OpenAI o1). Since these methods are based on different optimization dimensions, we can combine them feasibly in applications.
Specifically, some approaches prioritize a dominant paradigm for reasoning yet hope to integrate other paradigms for guidance. For instance, as shown in Figure 6 (b), the reasoning process of ToRA (Gou et al., 2024) presents an interleaved reasoning method with different knowledge media. It involves initial natural language generation, which is followed by code generation. Subsequently, the process awaits the results of code execution before generating further natural language. InternLM2.5-StepProver (Wu et al., 2024) exhibits a similar pattern. It incorporates natural language annotations to support symbolic reasoning steps. However, these interleaved reasoning approaches do not constitute true multi-paradigm reasoning. The primary paradigm can independently achieve the final answer without the supplementary paradigms. Consequently, these methods imply a single-paradigm reasoning approach enhanced by other paradigms, rather than genuine multi-paradigm reasoning.
Current methods often overlook the synergistic effects between paradigms and underestimate the importance of complete reasoning for achieving accurate results. To address this challenge, as shown in Figure 5 (c) and Figure 6 (c), we introduce Chain-of-Reasoning (CoR), a unified reasoning framework capable of multi-paradigm reasoning. This framework embodies the following potential advantages:
Sequential reasoning dependency. As a type of multi-paradigm reasoning, CoR is not a mere accumulation of isolated steps; instead, it represents a coherent, interconnected process. The output from an earlier paradigm functions not only as input for a subsequent paradigm but also informs the foundational information for its reasoning. For instance, knowledge derived from a natural language paradigm guides subsequent algorithmic paradigms, thereby enhancing the efficiency and accuracy of code generation based on detailed natural language reasoning. This mechanism also provides substantial context from prior paradigms to later paradigms and operates as a form of preliminary rehearsal where subsequent paradigms can follow correct reasoning steps or rectify incorrect ones.
A novel direction for test-time scaling emerges. We observe that the achievable improvements of single-paradigm reasoning are increasingly constrained by various search methods applied to the solution space. For instance, DeepSeek-Prover-V1.5-RL + RMaxTS requires $32\times 6400$ reasoning paths to achieve $63.5\%$ few-shot accuracy on the miniF2F benchmark. While extending the number of candidate solution paths seems a natural way to enhance the hit rate for ground truth solutions, the substantial computational effort involved underscores the inherent limitations of single-paradigm reasoning. This observation suggests that scaling up reasoning within a single paradigm during test time is inadequate for addressing complex tasks, such as mathematical problem solving. From a broader perspective, our CoR framework reexamines this challenge by shifting the focus of scaling efforts from reasoning paths to reasoning paradigms, thereby proposing a novel avenue for advancement.
Expanding the solution space. In multi-paradigm reasoning, the solution space is expanded by considering potential orthogonal relationships between reasoning paradigms. This allows generated solutions to be searched within different paradigms. Moreover, diverse solutions across paradigms can leverage chain relationships in reasoning to mutually inform each other. Therefore, our CoR framework enhances both intra-paradigm and inter-paradigm search capabilities within a large solution space. This diversity increases the likelihood of discovering optimal solutions.
Journey-based reasoning learning. CoR enhances models’ learning trajectory, allowing them to perform deep reasoning both within specific paradigms and across multiple paradigms. This approach expands the “journey” by introducing the ability to navigate and collaborate across diverse reasoning paradigms.
Compatibility with existing methods. Since most current approaches are fundamentally based on a single paradigm, CoR can seamlessly integrate these methods within its specified paradigms, enabling them to work synergistically and leverage their strengths. For example, CoT can be incorporated as a specific implementation of the NLR paradigm in CoR. This implies that CoR serves as a platform that facilitates continuous integration and collaboration.
7 Conclusions
This paper introduces Chain-of-Reasoning (CoR), a novel unified reasoning framework that synergistically combines natural language, algorithmic, and symbolic reasoning to solve mathematical problems. Our approach demonstrates the potential of integrating diverse reasoning paradigms in the reasoning process, enabling language models to tackle complex mathematical challenges effectively. CoR achieves significant performance gains on challenging mathematical reasoning tasks, substantially outperforming existing single-paradigm approaches on benchmarks like MATH and MiniF2F. In addition, CoR introduces a new, orthogonal perspective to traditional test-time scaling by leveraging multi-paradigm reasoning paths to enhance computational efficiency and performance. Furthermore, our proposed CoR framework offers valuable insights for future research, contributing to the development of more robust and versatile reasoning agents capable of leveraging the strengths of multiple reasoning modalities.
Appendix
Appendix A Experiment Details
A.1 The detail of the Universal Text Template
Universal Template Designed for Multi-paradigm Reasoning. ### Problem: [Statement of the problem] Let’s go through this step-by-step: 1. [Step 1 description] 2. [Step 2 description] 3. [Step 3 description] $\cdots$ [Further steps as needed] $\checkmark$ [Verification or conclusion of the step-by-step reasoning] Let’s write the corresponding formal proof in Lean 4 to prove this: ### Formal proof in Lean 4
[Lean 4 code for formal proof]
[Lean 4 Output]
Let’s use Python to perform these calculations: ### Code in Python
[Python code for calculation]
[Python Output]
### Summary [Summary of the solution and key takeaways]
Figure 7: The universal text template designed for multi-paradigm reasoning. The contents for each paradigm have been omitted.
As shown in Figure 7, we apply the designed universal text template on all training samples.
A.2 Example Prompts for Dataset Enhancement
To guide LLMs in generating and refining reasoning paradigms during dataset enhancement, we developed specific prompts ($p_{s}$) tailored to each seed dataset. These prompts provide structured instructions for the models to follow, ensuring the generation of high-quality and logically consistent reasoning paradigms.
Figure 9 and Figure 8 provide examples of such prompts designed for augmenting and refining formal proofs within the Lean 4 theorem prover for the Numina-TIR and Lean-Workbook datasets, respectively. The prompt specifies the input and output format and includes placeholders for the problem statement, informal proof, and corresponding formal proof in Lean 4. It also incorporates $N_{\text{shot}}$ few-shot examples to further guide the model. In our experiments, consistent with the common settings employed in (Jiang et al., 2023; Yang et al., 2024), we use $N_{\text{shot}}=8$. This structured approach and setting are adapted and applied to all datasets to ensure comprehensive coverage and logical consistency across the enhanced dataset.
Example Prompts for TIR Dataset Enhancement System Prompt:
You are an expert in Lean 4. Please respond to a math problem by translating the provided informal proof into Lean 4 code. Follow the format provided in the prompt. Please note that the informal proof and the formal proof need to be identical. Follow the format provided in the prompt.
User Input:
Now please translate the formal solution in Lean 4 following the instructions below. Please write the corresponding solution in Lean 4 (indicated by “Formal proof in Lean 4: ”) given the “# Problem: ” and “# Informal proof: ”, filling in the “# Formal proof in Lean 4: ” section.
You must respond in the following format: # Problem: … # Informal proof: … # Formal proof in Lean 4:
‘‘‘lean4
(lean 4 code for proving)
Here are examples you may refer to: $\mathrm{N}$ few-shot examples # Problem: {problem} # Informal proof: {informal_proof} # Formal proof in Lean 4:
‘‘‘lean4
Figure 8: Example prompts for Numina-TIR dataset enhancement. The few-shot examples provided for the LLM have been omitted.
Example Prompts for Lean-Workbook Dataset Enhancement System Prompt:
You are an expert in Lean 4 and formal mathematics. Your task is to explain how to construct formal proofs using Lean 4 syntax and tactics. Focus on the Lean 4 approach to theorem proving, rather than traditional mathematical reasoning.
User Input:
Please follow the instructions below to convert the Lean 4 code (indicated by “Formal proof in Lean 4: ”) into its informal proof, using the informal problem (indicated by “Problem: ”) as a guide. Please write the corresponding informal solution in natural language (indicated by “Informal proof: ”) given the “# Problem: ” and “# Formal proof in Lean 4: ”, filling in the “# Informal proof: ” section.
<Instruction> Analyze the given mathematical theorem and the corresponding Lean 4 code. Provide a detailed explanation of the proof process, adhering to the following guidelines: 1. Theorem structure: Clearly state the theorem, including its assumptions and conclusion. 2. Proof strategy: Explain the overall strategy employed in the proof, focusing on logical reasoning and mathematical deduction rather than calculation. 3. Step-by-step reasoning: Provide a detailed, step-by-step explanation of the proof process, ensuring that each step corresponds to an element in the Lean 4 code. 4. Logical deduction: Emphasize how each step of the proof follows logically from the previous steps or from the given assumptions. 5. Mathematical concepts: Discuss any specific mathematical concepts, notations, or definitions used in the proof, such as divisibility or exponentiation. 6. Abstraction: Present the proof in a general, abstract form that could be applied to similar problems, rather than focusing on specific numerical calculations. 7. Correspondence to code: Ensure that the logical flow of your proof explanation aligns with the structure and tactics used in the Lean 4 code, without explicitly mentioning Lean-specific terminology. Avoid using syntax or terminology specific to formal proof systems. Instead, focus on presenting a rigorous mathematical argument using general logical principles and mathematical language. The proof should follow the reasoning path implied by the Lean 4 code but be accessible to readers unfamiliar with formal proof assistants. </Instruction> You must respond in the following format: # Problem: … # Tags: … # Formal proof in Lean 4:
‘‘‘lean4
(lean 4 code for proving)
# Informal proof:
(Informal reasoning path for proving the problem)
Here are examples you may refer to: $\mathrm{N}$ few-shot examples # Problem: {problem} # Tags: {tag} # Formal proof in Lean 4:
‘‘‘lean4
{lean_workbook code}
# Informal proof: …
Figure 9: Example pompts for Lean-Workbook dataset enhancement. The few-shot examples provided for the LLM have been omitted.
A.3 Details of Training Settings
In all stages of the PPT method, a learning rate of $52e-$ 5 and a warm-up ratio of $1\%$ are implemented. To enhance computational efficiency, the training process is conducted using distributed optimization with DeepSpeed ZeRO Stage 3 (Rajbhandari et al., 2020) and is combined with Flash-Attention (Dao, 2024). Stage ➀ comprises $3$ epochs, whereas the satge ➁ and ➂ are conducted over $4$ epochs. A sequence length of $2,048$ tokens is used in Stages ➀ and ➁, which is increased to $4,096$ tokens in Stage ➂ to support more complex reasoning tasks. Furthermore, we employ an annealing strategy at the end of Stage ➂ with high-quality MPM samples, with the aim of further enhancing model accuracy in complex reasoning tasks.
A.4 Benchmarks
We report the statistics of the evaluation datasets in Table 6.
Table 6: The statistics of the evaluation datasets.
Dataset # of Test set Avg. Length
MATH 5,000 30.7
GSM8K 1,319 46.3
MiniF2f 244 30.5
A.5 Details of Metrics
This study employs the pass@$N$ metric on the MiniF2F benchmark, with the evaluation grounded in a sample budget of $N$. To ensure a fair comparison of computational cost across different generation schemes, this paper defines the sample budget according to the following rules.
  • For single-pass sampling methods, we define the sample budget $N$ as the total number of proofs generated, with large values of $N$ factorized for ease of comparison to tree search methods.
  • For best-first-search methods, following the notation in Llemma (Azerbayev et al., 2024), we present $N=N_{\text{att}}\times N_{\text{tact}}\times N_{\text{iter}}$ where $N_{\text{att}}$ denotes the number of best-first-search attempts, $N_{\text{tact}}$ denotes the number of tactics generated for each expansion, and $N_{\text{iter}}$ denotes the number of expansion iterations.
  • For tree-based search methods, e.g., RMaxTS (Xin et al., 2024) and HTPS (Lample et al., 2022), we present $N=N_{\text{att}}\times N_{\text{ex}}$, where $N_{\text{att}}$ denotes the number of tree search attempts, and $N_{\text{ex}}$ denotes the number of model generations invoked in tree expansions.
  • For our SMPS, we define the sample budget as $N=N_{\text{NLR}}\times N_{\text{SR}}$, where $N_{\mathrm{NLR}}$ denotes the number of initial semantic reasoning paths, and $N_{\mathrm{SR}}$ denotes the number of symbolic reasoning paths extended for each other reasoning paradigm.
Appendix B Additional Analysis
B.1 The Risk of Data Leakage
We measure the Levenshtein distance (Miller et al., 2009) between problems in the training dataset and those in the test dataset to mitigate the risk of data leakage. During the experiments, we set a similarity threshold of $0.7$ and excluded cases from the training dataset that exceeded this threshold. As a result, approximately $1,000$ cases, accounting for less than $0.6\%$ of our training dataset, were removed to prevent potential leakage.
B.2 Different Evaluation Strategies on Arithmetic Benchmarks
Table 7: Zero-shot results of different evaluation strategies across MATH and GSM8k benchmarks.
Model MATH GSM8k
Pass@1 Maj@8 Pass@1 Maj@8
Llama-3.1-8B 58.18 59.46 84.00 87.26
DeepSeek-Math-Base-7B 66.74 71.70 88.70 91.43
Table 7 explores the impact of various evaluation strategies on the CoR framework. The majority vote strategy benefits the CoR framework. Specifically, a majority vote strategy with $8$ candidate samples on GSM8K can exceed GPT-4o’s Pass@1 performance ($90.5\%$).

### v4

@@
 We propose a Progressive Paradigm Training (PPT) strategy that allows models to progressively master these paradigms, culminating in the development of CoR-Math-7B.
-Experimental results demonstrate that CoR-Math-7B significantly outperforms current SOTA models, achieving up to a $41.0\%$ absolute improvement over GPT-4 in theorem proving tasks and a $7.9\%$ improvement over RL-based methods in arithmetic tasks.
-These results showcase the enhanced mathematical comprehensive ability of our model, achieving significant performance gains on specific tasks and enabling zero-shot generalization across tasks.
-Figure 1: A comprehensive comparative analysis of CoR-Math-7B and baseline models across challenging mathematical tasks.
-(a) demonstrates the effectiveness of CoR-Math-7B (zero-shot) in theorem proving tasks, specifically emphasizing its advantage over baseline and their fine-tuned variants over SFT and RL.
-(b) shows a resource-efficiency analysis for arithmetic computation tasks, plotting model performance vs.
-SFT data size, in which CoR-Math-7B attains the optimal resource efficiency, achieving near-optimal zero-shot performance.
+Experimental results demonstrate that CoR-Math-7B significantly outperforms current SOTA models, achieving up to a $41.0\%$ absolute improvement over GPT-4o in theorem proving tasks and a $15\%$ improvement over RL-based methods on the MATH benchmark in arithmetic tasks.
+These results show the enhanced mathematical comprehensive ability of our model, enabling zero-shot generalization across tasks.
+The code is available at https://github.com/microsoft/CoR.
+[…]
+[…]
+[Uncaptioned image] Chain-of-Reasoning: Towards Unified Mathematical Reasoning in Large Language Models via a Multi-Paradigm Perspective
+[…]
 1 Introduction
-While LLMs have demonstrated strong performance in solving mathematical tasks (Feigenbaum et al., 1963; Hosseini et al., 2014), advanced open-source reasoners still struggle with solving comprehensive mathematical problems, including both arithmetic computation and theorem proving.
+Figure 1: A comprehensive comparative analysis of CoR-Math-7B and baseline models across mathematical tasks.
+(a) shows the effectiveness of CoR-Math-7B (zero-shot) in theorem proving tasks.
+(b) shows a resource-efficiency analysis for arithmetic computation tasks, where CoR-Math-7B achieves optimal resource efficiency and near-optimal zero-shot performance.
+While LLMs have shown strong performance in solving mathematical tasks (Feigenbaum et al., 1963; Hosseini et al., 2014), advanced open-source reasoners still struggle with solving comprehensive mathematical problems, including both arithmetic computation and theorem proving.
+Existing works (Xin et al., 2024; Yang et al., 2024; Wu et al., 2024; Zhang et al., 2024) are often trained on specific tasks, aiming to enhance their ability to independently derive answers based on specific structured knowledge representation.
+This representation is known as the reasoning paradigm, involving Natural Language Reasoning (NLR), Algorithmic Reasoning (AR), and Symbolic Reasoning (SR), as depicted in Fig.˜2 (a).
+Specifically, NLR uses natural language for reasoning via common sense and semantic context, providing explicit step-by-step explanations (Wei et al., 2022).
+AR leverages code to emphasize computational operations and execution processes, such as generating Python code for execution (Chen et al., 2023; Gao et al., 2023).
+SR uses logical symbols and axiomatic systems for formalized reasoning, with recent methods (Xin et al., 2024; Huang et al., 2024; Wu et al., 2024) considering numerous symbolic trajectories via tree-based search for theorem proving.
+However, these methods mainly focus on optimizing single-paradigm reasoning, creating models that demonstrate asymmetrical performance across different mathematical tasks.
+For instance, models specialized in NLR may exhibit deficiencies in theorem proving and vice versa.
+This fragmented paradigm constrains the upper-bound performance on individual tasks and undermines the model’s capacity for cross-paradigm generalization.
 Figure 2: The reasoning process under different paradigms: (a) In single-paradigm reasoning, each reasoning step relies on the same knowledge medium, such as Natural Language (NL), algorithms, or symbols.
@@
 (c) The proposed CoR reasoning framework, along with several examples, shows that multi-paradigm reasoning allows for varying reasoning depths to address different types of problems.
-Existing works (Xin et al., 2024; Yang et al., 2024; Wu et al., 2024) are often trained on specific tasks, aiming to enhance their ability to independently derive answer based on specific structured knowledge representation.
-This representation is known as the reasoning paradigm, involving Natural Language Reasoning (NLR), Algorithmic Reasoning (AR), and Symbolic Reasoning (SR), as illustrated in Figure 2 (a).
-Specifically, NLR leverages natural language text for reasoning based on human common sense and semantic context, with explicit step-by-step natural language explanations (Wei et al., 2022).
-AR leverages code to focus on the computer’s operations and execution process, performing inference on the final target, such as the generation of Python code for compilation (Chen et al., 2023; Gao et al., 2023) to obtain the final answer.
-SR utilizes logical symbols and axiomatic systems as the fundamental tools for rigorously formalized reasoning, with current methods (Xin et al., 2024; Huang et al., 2024; Wu et al., 2024) considering numerous symbolic trajectories through a tree-based search process for theorem proving.
-However, these methods focus on improving single-paradigm reasoning and overlook the potential of collaboration among multiple paradigms, which restricts their single-task performance and hinders their cross-task generalization.
 Researchers have explored various strategies to tackle these challenges.
-To improve the single-task performance, some works have explored integrating tools to overcome the limitations of single-paradigm reasoning (Gou et al., 2024; LI et al., 2024), as shown in Figure 2 (b).
-These methods combine natural language with code-based algorithms (tools) to enable interleaved reasoning and generate a final answer.
-Although these methods acknowledge the potential benefits of integrating different reasoning paradigms, they often rely on a single paradigm to complete the reasoning process.
-They overlook the possibility that the second paradigm could independently complete the reasoning, thereby constraining its overall potential.
-Besides, in an effort to overcome the challenge of achieving cross-task generalization, several studies (Shao et al., 2024; Huang et al., 2024) incorporate data samples from diverse task types into large-scale training datasets, such as those drawn from theorem proving tasks that focus exclusively on SR solutions, or from arithmetic problems that emphasize NLR solutions.
+To improve the single-task performance, some works integrate tools to overcome the limitations of single-paradigm reasoning (Gou et al., 2024; LI et al., 2024), as shown in Fig.˜2 (b).
+While recognizing the benefits of combining reasoning paradigms, they still rely on a single paradigm, neglecting the possibility that the second paradigm could independently complete the reasoning, thereby constraining overall potential.
+Besides, to improve cross-task generalization, several studies (Shao et al., 2024; Huang et al., 2024) incorporate diverse task samples into large-scale datasets, such as those drawn from theorem proving tasks that focus exclusively on SR solutions, or from arithmetic problems that emphasize NLR solutions.
 Although models trained on such data are capable of cross-task reasoning, they still rely on demonstrations for effective transfer.
-To address these limitations, we propose a novel unified reasoning framework, Chain-of-Reasoning (CoR), which chains NLR, AR, and SR together to generate synergistic benefits.
-As illustrated in Figure 2 (c), CoR enables multi-paradigm reasoning for a given problem, which applies different reasoning paradigms to derive multiple potential answers and are then summarized into a final answer.
-The framework allows the model to continue reasoning using additional paradigms based on previously generated ones, facilitating collaboration among paradigms to enhance individual task performance.
-Moreover, CoR implements unified multi-paradigm reasoning across different tasks to obtain the required answers, thereby achieving zero-shot reasoning across tasks.
-In detail, by adjusting the prompts, the depth of reasoning can be varied, which improves the model’s adaptability to diverse tasks.
-By constructing such multi-paradigm reasoning paths, we create a training dataset named MPM, comprising $167$k reasoning paths.
-Furthermore, we propose a novel Progressive Paradigm Training (PPT) approach, which enables models to progressively master multiple reasoning paradigms, resulting in CoR-Math-7B.
-We evaluated CoR-Math-7B on three representative and challenging mathematical reasoning benchmarks that effectively demonstrate comprehensive reasoning capabilities, encompassing both arithmetic computation and theorem proving.
-These benchmarks include MATH (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), and MiniF2F (Zheng et al., 2022).
-Our results show that LLMs equipped with CoR significantly outperform current SOTA baselines.
-In the theorem proving task as shown in Figure 1 (a), CoR-Math-7B surpasses GPT-4o’s (Hurst et al., 2024) zero-shot performance by a remarkable absolute $41.0\%$ and outperforms all other few-shot reasoners.
-In arithmetic tasks, our CoR-Math-7B first performs NLR and SR as a rehearsal, which enhances the accuracy of AR and demonstrates the effectiveness of synergy.
-For example, CoR-Math-7B achieves an absolute improvement of $7.9\%$ over Reinforcement Learning (RL) methods (Shao et al., 2024) that utilize the same base model on the MATH dataset.
-Furthermore, as shown in Figure 1 (b), CoR-Math-7B efficiently utilizes resources to surpass the optimal performance curve of single-paradigm approaches.
-Compared to mainstream methods that perform extensive searches within a single paradigm, our CoR framework achieves better results by extending test-time inference across multiple paradigm spaces.
-These results show that our method achieves comprehensive mathematical problem solving abilities though multi-paradigm reasoning, requiring fewer training data and reducing reasoning costs.
+To address these limitations, we propose Chain-of-Reasoning (CoR), a unified framework integrating NLR, AR, and SR to produce synergistic benefits.
+As shown in Fig.˜2 (c), CoR enables multi-paradigm reasoning by applying different reasoning paradigms to derive multiple potential answers, which are then summarized into a final solution.
+This framework allows iterative reasoning across paradigms, leveraging prior results to enhance single-task performance.
+Moreover, CoR unifies multi-paradigm reasoning across tasks, enabling zero-shot generalization.
+Adjusting prompts to control reasoning depth improves adaptability to diverse tasks.
+As a result, we introduce Multi-Paradigm Mathematical (MPM), a dataset comprising 167167k reasoning paths, and propose Progressive Paradigm Training (PPT), a method enabling models to progressively master multiple reasoning paradigms, leading to CoR-Math-7B.
+We evaluate CoR-Math-7B on five challenging mathematical reasoning benchmarks, covering both arithmetic computation and theorem proving.
+Our results show that LLMs equipped with CoR significantly surpass current SOTA baselines.
+In theorem proving (Fig.˜1 (a)), CoR-Math-7B improves zero-shot performance over GPT-4o (Hurst et al., 2024) by $41.0\%$, outperforming all few-shot reasoners.
+For arithmetic tasks, CoR-Math-7B achieves a $24.2\%$ absolute improvement over GPT-4 on MATH.
+Furthermore, as shown in Fig.˜1 (b), CoR-Math-7B efficiently utilizes resources to surpass the optimal performance curve of single-paradigm approaches.
+Unlike mainstream methods that conduct extensive searches within a single paradigm, CoR enhances test-time inference in multiple paradigms.
+These results show that CoR can solve comprehensive mathematical problems through multi-paradigm reasoning, requiring less training data and lowering reasoning costs.
 2 Related Work
 Reasoning Paradigms in LLMs.
-Recent advancements in LLMs have primarily concentrated on enhancing single-paradigm reasoning, where each paradigm serves as a distinct medium for knowledge representation and logical inference.
-NLR utilizes human-readable language to solve commonsense reasoning and step-by-step deductions (Wei et al., 2022; Yao et al., 2023; Zhou et al., 2023; Besta et al., 2024; Sel et al., 2024).
-For precise numerical computations, AR is to generate executable code, such as Python scripts, to perform accurate calculations through execution (Chen et al., 2023; Rozière et al., 2023).
-In the realm of theorem proving, SR leverages logical symbols and axiomatic systems to ensure rigorous and verifiable proofs (Xin et al., 2024).
-Furthermore, some approaches leverage external tools within the single-paradigm to enhance reasoning capabilities (Gou et al., 2024), such as using calculators or code interpreters, while maintaining the core reasoning process within a single paradigm.
-While these paradigms excel within their specific areas, they often struggle with cross-domain generalization and dynamic environments.
-To address these issues, the CoR framework facilitates collaboration between various reasoning paradigms to enable zero-shot multitasking generalization.
+Recent advancements in LLMs focus on single-paradigm reasoning, with each paradigm representing a distinct method for knowledge representation and logical inference.
+NLR uses human-readable language for commonsense reasoning and step-by-step deductions (Wei et al., 2022; Yao et al., 2023; Zhou et al., 2023; Besta et al., 2024; Sel et al., 2024), AR generates executable code for precise calculations (Chen et al., 2023; Rozière et al., 2023), and SR utilizes logical symbols for theorem proving (Xin et al., 2024).
+Some methods enhance reasoning by integrating external tools like calculators or code interpreters, yet they remain within a single paradigm (Gou et al., 2024).
+While effective in specific tasks, they struggle with cross-domain generalization and dynamic environments.
+To address these limitations, the CoR framework enables collaboration across reasoning paradigms, enabling zero-shot multitask generalization.
 Mathematical Problem Solving with LLMs.
-Advanced research demonstrates the significant potential of LLMs in solving mathematical problems (Zhu et al., 2023).
-Several studies have developed unified mathematical solvers by synthesizing mathematical data, capable of addressing various challenges such as arithmetic calculations and theorem proving (Huang et al., 2024; Shao et al., 2024).
-However, recent research tends to focus on optimizing single paradigms.
-For instance, in the field of arithmetic computation, there is a tendency to incorporate programming tools to aid natural language mathematical reasoning (Gou et al., 2024).
-In theorem proving, researchers not only utilize specialized annotated data, including pre-training data and reinforcement learning reward data, but also employ tree-based search methods to generate numerous possible solutions (Ying et al., 2024b; Xin et al., 2024).
-These methods either fail to fully consider the importance of complete reasoning or rely heavily on large-scale search within the solution space to fully explore the performance improvement potential of a single paradigm.
-To address these issues, we introduce the MPM dataset and present the CoR-Math-7B model, trained on it, which leverages the synergistic effects of multi-paradigm reasoning to navigate towards the optimal solution within the expanded multi-paradigm solution space.
+Advanced research highlights the potential of LLMs in solving mathematical problems (Zhu et al., 2023; Lin et al., 2024b; Luo et al., 2025).
+Several studies have developed unified solvers by synthesizing mathematical data to address challenges like arithmetic computation and theorem proving (Huang et al., 2024; Shao et al., 2024).
+However, most approaches focus on optimizing a single paradigm.
+For instance, in arithmetic computation, tools are integrated to assist natural language reasoning.
+(Gou et al., 2024).
+In theorem proving, models rely on specialized data like pre-training data and Reinforcement Learning (RL) reward data, and tree-based search methods to generate numerous possible solutions (Ying et al., 2024b; Xin et al., 2024).
+These approaches either neglect complete reasoning or depend on large-scale search within the solution space, limiting performance gains within a single paradigm.
+To address these, we introduce the CoR-Math-7B model with complete reasoning processes to explore an expanded multi-paradigm solution space.
 Figure 3: An overview of (a) the Multi-Paradigm Math (MPM) dataset construction process, involving reconstruction, extension, and theorem prover verification, and (b) the Progressive Paradigm Training (PPT) method, where the model is trained with increasing reasoning paradigms in stages.
@@
 3.1 Overview
-Our CoR framework aims to enable LLMs to perform a series of multi-paradigm reasoning on any type of mathematical problem, ultimately arriving at a solution.
+CoR aims to enable LLMs to perform a series of multi-paradigm reasoning on any type of mathematical problem, ultimately arriving at a solution.
 Specifically, given a mathematical problem $x$, LLMs ($\mathbb{P}$) can infer the result $y$ by following multiple reasoning paradigms, where each reasoning paradigm $\tau$ includes $n$ reasoning paths $\{rp_{1},...,rp_{n}\}$.
-To formally represent this single-paradigm scenario, we can formulate it as $y\sim\mathbb{P}(y|x,\tau)$.
-To simplify the process for each reasoning paradigm, we set $n=1$ as default (Details in Section 6).
+We represent this single-paradigm scenario as $y\sim\mathbb{P}(y|x,\tau)$.
+To simplify the process for each reasoning paradigm, we set $n=1$ as default (Details in Appendix˜A).
 Inspired by recent works (Wei et al., 2022) in encouraging step-by-step reasoning, we introduce CoR, which extends from a single paradigm to three paradigms $\Gamma=(\tau_{1},\tau_{2},\tau_{3})$.
 For generating multiple chained reasoning paradigms for a given problem, CoR follows the steps outlined below.
-The reasoning process begins with the initialization of the problem $x$ and the first reasoning paradigm $\tau_{1}$.
+The reasoning process begins with the problem $x$ and the first reasoning paradigm $\tau_{1}$.
 Subsequently, each paradigm $\tau_{i}$ in the sequence $\Gamma$ is generated based on the problem $x$ and the previously established paradigms $\tau_{1},...,\tau_{i-1}$, as represented by $\tau_{i}\sim\mathbb{P}(\tau_{i}|x,\tau_{1},...,\tau_{i-1})$.
-Finally, the outcomes from all reasoning paradigms are aggregated to derive the final result $y$, expressed as $y\sim\mathbb{P}(y|x,\tau_{\text{NLR}},\tau_{\text{SR}},\tau_{\text{AR}})$, considering three paradigms: NLR, SR, and AR.
+Finally, the outcomes of all the reasoning paradigms are aggregated to derive the final result $y$, expressed as $y\sim\mathbb{P}(y|x,\tau_{\text{NLR}},\tau_{\text{SR}},\tau_{\text{AR}})$, considering three paradigms: NLR, SR, and AR.
 In detail, $\tau_{\text{SR}}$ and $\tau_{\text{AR}}$ include both the complete reasoning paths and the interaction processes with tool outputs.
-In our experiments, $\tau_{\text{SR}}$ utilizes the Lean prover, and $\tau_{\text{AR}}$ leverages a Python compiler as the algorithmic interactive tool.
-CoR possesses the capacity for broad application in complex reasoning scenarios.
-In the context of mathematical reasoning, it manifests as CoR-Math, with potential for future adaptations such as CoR-Code.
-According to this framework, as shown in Figure 3, the training pipeline consists of three key phases: (a) collecting a Multi-Paradigm Mathematical (MPM) dataset, with deep reasoning paths based on multiple reasoning paradigms; (b) introducing a Progressive Paradigm Training (PPT) strategy to shape the model’s reasoning behavior, gradually mastering more reasoning paradigms; and (c) leveraging the trained LLM for zero-shot inference, allows multi-paradigm reasoning with adaptable depth, and uses Sequential Multi-Paradigm Sampling (SMPS) to explore a diverse solution space.
+In this study, $\tau_{\text{SR}}$ uses the Lean prover, and $\tau_{\text{AR}}$ applies a Python compiler to obtain reasoning results.
+In our workflow (Fig.˜3), the training pipeline includes: (a) collecting the MPM dataset with deep reasoning paths; (b) using progressive paradigm training to enhance reasoning across paradigms; and (c) applying the trained LLMs for zero-shot inference with sequential multi-paradigm sampling to explore diverse solutions.
 3.2 Collecting Dataset
-[…]
+To train CoR model, we extend the single-paradigm datasets to incorporate multiple reasoning paradigms, denoted as $\texttt{x},\texttt{NLR},\texttt{SR},\texttt{AR},\texttt{y}>$.
+As presented in Fig.˜3 (a), the training data collection process involves two stages: (a.1) Reconstructing and Extending, and (a.2) Revising.
+In the first stage, we reconstruct high-quality open-source mathematical data as seed samples and synthesize additional reasoning paradigms to form the MPM-raw dataset.
+In the second stage, a theorem prover examines MPM-raw samples, and a mathematical reasoner revises those that fail, compiling all completed data into the final MPM dataset.
 Stage 1: Reconstructing and Extending.
-We introduce a universal text template designed for multi-paradigm reasoning (Details in Section A.1).
-This template standardizes the positioning of distinct reasoning paradigms and defines the relationships between them.
-It accommodates various reasoning depths and supports flexible combinations of different reasoning paradigms.
-As shown in Figure 2 (c), Problem 1 demonstrates an instance incorporating both NLR and SR paradigms, while Problem 2 showcases an instance that integrates three reasoning paradigms.
-To ensure data integrity and prevent potential biases, we pre-process the Numina-TIR (LI et al., 2024) and Lean-Workbook (Ying et al., 2024a) datasets through a two-step process.
+We introduce a universal text template for multi-paradigm reasoning (details in Sec.˜B.1), which standardizes the placement and relationships of reasoning paradigms.
+It supports various reasoning depths and flexible combinations of different reasoning paradigms.
+As shown in Fig.˜2 (c), Problem 1 shows an instance incorporating both NLR and SR paradigms, and Problem 2 integrates three reasoning paradigms.
+To ensure data integrity and prevent potential biases, we pre-process the Numina-TIR (LI et al., 2024) and Lean-Workbook (Ying et al., 2024a) datasets in two steps.
 First, samples without corresponding solutions are removed.
@@
 These models generate missing reasoning paradigms $\tau_{g}$ and refine existing ones $\tau^{\prime}$, ensuring comprehensive coverage and logical consistency.
-To effectively guide the processes of augmentation and refinement, we develop tailored prompts $p_{s}$ for each seed dataset.
-Examples of these prompts are detailed in Section A.2.
+To effectively guide the processes of augmentation and refinement, we develop tailored prompts $p_{s}$ for each seed dataset (examples in Sec.˜B.2).
 The procedure can be described in:
@@
 This approach considerably lowers the requisite skill level of annotators.
-Furthermore, to avoid data leakage, we calculate the Levenshtein distance (Miller et al., 2009) between problems in the training and test samples, while details are provided in the Section B.1.
+To prevent data leakage, we compute the Levenshtein distance (Miller et al., 2009) between training and test problems, as detailed in Sec.˜C.1.
 As a result, this phase yields the MPM-raw dataset, comprising approximately $285,000$ synthetic samples.
@@
 The MPM-raw dataset interacts with the Lean prover to verify proof steps, guiding the filtering and modification of reasoning paths.
-The process submits the proof $\tau_{\text{SR}}$ to the prover.
-Once the prover successfully completes the proof without returning any errors, the entire multi-paradigm reasoning path is collected to the MPM dataset.
+The proof $\tau_{\text{SR}}$ is submitted to the prover, and if verification is successful without errors, the entire multi-paradigm reasoning path is collected to the MPM dataset.
 Otherwise, the error information $\varepsilon$ returned by the prover is fed into a revising model $\mathbb{P}_{R}$.
@@
 where the revised proof $\tilde{\tau}_{\text{SR}}$ is then resubmitted to the Lean prover for verification.
-This iterative process continues for a maximum of $64$ iterations or until the prover verifies $\tilde{\tau}_{\text{SR}}$ as correct.
-In detail, we utilize DeepSeek-Prover-V1.5 (Xin et al., 2024) as the revising model.
-As a result, the MPM dataset comprises $82,770$ problems and $167,412$ multi-paradigm reasoning solutions.
+This iterative process continues for up to 64 iterations or until $\tilde{\tau}_{\text{SR}}$ is verified as correct.
+In detail, we employ DeepSeek-Prover-V1.5 (Xin et al., 2024) as the $\mathbb{P}_{R}$.
+Consequently, the MPM dataset comprises $82,770$ problems and $167,412$ multi-paradigm reasoning solutions.
 3.3 Training
-Inspired by recent advances (LI et al., 2024), we introduce the Progressive Paradigm Training (PPT) strategy, which enables LLMs to gradually master an increasing variety of reasoning paradigms.
-As shown in Figure 3 (b), this training framework enables the model to progressively expand its reasoning capability by introducing different types of reasoning data in stages.
+Inspired by recent advances (LI et al., 2024), we introduce the Progressive Paradigm Training (PPT) strategy, enabling LLMs to gradually master diverse reasoning paradigms.
+As shown in Fig.˜3 (b), this framework expands the model’s reasoning abilities by sequentially introducing different paradigms at each stage.
 Each training stage uses a different combination of reasoning paradigms.
-In the stage ➀, given the dominance of NL in the pre-training data of language models (Dubey et al., 2024), we create Numina-CoT as an initialized teaching stage.
-This is done by modifying the original Numina-CoT dataset (LI et al., 2024) according to our universal text template, which enables the model to learn to use NL to solve complex mathematical problems.
+In stage ➀, given the dominance of NL in the pre-training data of language models (Dubey et al., 2024), we create Numina-CoT as an initialized teaching stage.
+This is done by modifying the original Numina-CoT dataset (LI et al., 2024) according to our universal text template, enabling the model to learn to use NL to solve complex mathematical problems.
 Based on the question $x$, the model performs reasoning $\tau_{\text{NLR}}$ and generates the answer $y$.
-The generated sequence is $z=[x]\tau_{\text{NLR}}y$, where $[x]$ represents the tokens of the inputs.
+The generated sequence is $z=[x]\tau_{\text{NLR}}y$, where $[x]$ represents the inputs after tokenization.
 For simplicity, we first consider the loss function for a single sample:
@@
 After this stage, the model can handle problems that require precise answers.
-Finally, in the stage ➂, we further expand the training data to three reasoning paradigms by utilizing the MPM dataset, where $z$ is $[x]\tau_{\text{NLR}}\tau_{\text{AR}}\tau_{\text{SR}}y$.
-After full PPT stages, the trained CoR-Math-7B model not only masters NLR and AR but also can perform rigorous logical SR.
+In stage ➂, we further expand the training data to three reasoning paradigms by utilizing the MPM dataset, where $z$ is $x]\tau_{\text{NLR}}\tau_{\text{AR}}\tau_{\text{SR}}y$.
+After full stages, the trained CoR-Math-7B model not only masters NLR and AR but also can perform rigorous logical SR.
+Unlike traditional curriculum and incremental learning, which sequencing tasks by increasing difficulty within a single paradigm or accumulating knowledge, PPT progressively integrates fundamentally distinct reasoning paradigms, transitioning from familiar to unfamiliar, while fostering synergy across paradigms.
 3.4 Inference
-We introduce an inference method that combines variable reasoning depth and sequential multi-paradigm sampling to solve comprehensive mathematical tasks, enabling the model to adapt its reasoning process based on task-specific requirements.
+We present an inference method combining variable reasoning depth and multi-paradigm sampling to solve comprehensive mathematical tasks.
 Prompts with variable reasoning depth.
-In the zero-shot inference phase, the CoR-Math-7B exhibits proficiency in multi-paradigm reasoning.
-This proficiency allows for the implementation of multi-paradigm reasoning with adaptable reasoning depths, achieved by adjusting prompts according to the specific demands of a given task.
+In the zero-shot inference phase, CoR-Math-7B exhibits proficiency in multi-paradigm reasoning, enabling adjustable reasoning depths by modifying prompts based on task requirements.
 Initially, the model is prompted to conduct NLR, thereby activating a wide range of knowledge patterns associated with natural language from the pre-training corpus.
 Subsequently, the inference method is tailored to the specific problem type.
-For instance, in the context of theorem proving, the model transitions to SR, which is better suited for formal deduction.
-Given that the output of the model during SR is structured in Lean 4, the relevant proof segment can be extracted as the final solution.
-In cases of arithmetic computation, the model first performs NLR, followed by SR to strengthen logical coherence, and ultimately employs AR for precise calculations.
+For example, in theorem proving, the model switches to SR, which is more suitable for formal deduction.
+Since the SR output is structured in Lean 4, the relevant proof segment can be extracted as the final solution.
+In arithmetic computations, the model first employs NLR, followed by SR for logical coherence, and concludes with AR for precise calculations.
 A summary box is utilized to present the final results.
+The paradigm-switching behavior depends on the presence of prompts and can be categorized into instruction-followed reasoning and instruction-free reasoning.
+This section primarily presents results from the former.
+Instruction-free reasoning, which explores how the model autonomously switches paradigms without prompt guidance, will be demonstrated in  Sec.˜D.1.
 This adaptable paradigm effectively captures the distinct reasoning patterns inherent in different paradigms and demonstrates flexibility in accommodating a variety of scenarios.
@@
 Instead of token-level sampling based on tree structures (Qiu et al., 2024; Xiong et al., 2024) within single-paradigm reasoning paths, we purpose sequential paradigm-level sampling method, named SMPS.
-This approach allows the model to generate outputs by sequentially sampling across different paradigms.
+This allows the model to generate outputs by sampling across different paradigms.
 For instance, in a two-paradigm reasoning scenario, the model first instantiates $J$ distinct paths for the initial reasoning paradigm $\tau_{1}$.
 $\tau_{1j}\sim\mathbb{P}(\tau_{1j}\mid x),\quad\forall j\in{1,\dots,J}$ (4)
-Subsequently, for each of these $\tau_{1}$ paths, the model instantiates $K$ possible paths for the secondary reasoning paradigm $\tau_{2}$
+Subsequently, for each of these $\tau_{1}$ paths, the model instantiates $K$ possible paths for the secondary reasoning paradigm $\tau_{2}$.
 $\tau_{2k}\sim\mathbb{P}(\tau_{2k}\mid x,\tau_{1j}),\quad\forall k\in{1,\dots,K},\forall j$ (5)
@@
 $y_{jk}\sim\mathbb{P}(y_{jk}\mid x,\tau_{1j},\tau_{2k}),\quad\forall j,k$ (6)
-In summary, the SMPS method leverages the combinatorial expansion of reasoning paths to explore a diverse paradigm-based solution space comprehensively.
-This structured approach enhances the robustness and depth of the reasoning results.
+Overall, the SMPS method utilizes a combinatorial expansion of reasoning paths to explore a diverse paradigm-based solution space, enhancing the robustness and depth of the reasoning results.
 4 Experimental Settings
@@
 Datasets.
-To evaluate our model’s mathematical reasoning capabilities comprehensively, we utilize a suite of widely recognized and representative benchmark datasets.
-These datasets encompass a diverse range of mathematical reasoning tasks, including arithmetic computation and theorem proving.
-The arithmetic computation ability is evaluated on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) datasets.
-These datasets span a spectrum of difficulty levels, from elementary arithmetic to mathematical competitions.
+We conduct extensive experiments to evaluate the model’s mathematical reasoning abilities, including arithmetic computation and theorem proving.
+The arithmetic computation ability is evaluated on the GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), AMC2023 (AI-MO, 2024b) and AIME2024 (AI-MO, 2024a) datasets.
 Theorem proving ability is evaluated on the miniF2F test set (Zheng et al., 2022), which features mathematical problems of Olympiad difficulty.
-Metrics.
-Accuracy serves as the primary metric for performance evaluation across all datasets.
-In detail, we follow the widely-used CoT settings (Wei et al., 2022) to obtain solutions for arithmetic problem tasks, with numerical answers rounded to the nearest integer.
-The SymPy library [^1] https://www.sympy.org/ is utilized for parsing and evaluating mathematical expressions.
-To address potential variations in numerical representation, the model is instructed to explicitly state the final answer in the conclusion, following the approach in Numina models (LI et al., 2024).
-For theorem proving datasets, we follow the recent advances (Xin et al., 2024) to adapt the publicly available MiniF2F benchmark for the Lean 3 theorem proving environment to the Lean 4 environment [^2] https://github.com/rahul13613/miniF2F-lean4.
-These adjustments ensure the use of the most up-to-date formal verification tools.
-The pass@$N$ metric evaluates the capacity of the model to generate correct proofs within $N$ attempts.
-Specifically, our CoR-Math-7B employs SMPS, which operates by performing NLR followed by SR.
-Therefore, the sample budget $N$ is $N_{\text{NLR}}\times N_{\text{SR}}$.
-Detailed settings of all models are provided in Section A.5.
-4.2 Implementation Details
-We fine-tuned widely-used DeepSeek-Math-Base 7B (Shao et al., 2024) and Llama-3.1 8B (Dubey et al., 2024) models, employing our PPT method on MPM dataset.
-Unless otherwise specified, CoR-Math-7B model is based on DeepSeek-Math-Base 7B.
-The details are available in Section A.3.
-4.3 Baselines
-This study examines three categories of baseline models, with results reported using CoT prompting.
-General-purpose mathematical models: Considering the general applicability of our CoR-Math-7B model, we present the SOTA general-purpose mathematical models, including Mustard (Huang et al., 2024), DeepSeek-Math (Shao et al., 2024), InternLM-Math (Ying et al., 2024b), Llama-3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Llemma (Azerbayev et al., 2024).
-Task-specific mathematical models: We consider several expert models on mathematical optimization for specific tasks.
-For expert models on arithmetic computation, we include several models that undergo specialized reinforcement in arithmetic computation to achieve enhanced performance, such as more data and tool invocation.
-These models encompass Qwen2.5-Math (Yang et al., 2024), WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2024), DART-Math (Tong et al., 2024), InternLM-Math (Ying et al., 2024b), , DeepSeek-Math-Instruct / RL (Shao et al., 2024), Xwin-Math (Li et al., 2024), ToRA (Gou et al., 2024), and NuminaMath (LI et al., 2024).
-For expert models on theorem proving, we consider several SOTA models with substantial improvements, including reinforcement learning and Monte Carlo search.
-We include LLM-Step (Welleck and Saha, 2023), GPT-f (Polu and Sutskever, 2020), Lean-STaR (Lin et al., 2024), Hypertree Proof Search (Lample et al., 2022), DeepSeek-Prover (Xin et al., 2024), and InternLM2.5-StepProver (Wu et al., 2024).
-Foundation models and proprietary models: We introduce current advanced open-source foundational models, which are Llama-3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Mixtral (Jiang et al., 2024).
-We also introduce proprietary models, which are GPT-4 (OpenAI, 2023), GPT-4o (Hurst et al., 2024) and o1-mini (OpenAI, 2024), as robust competitors.
-5 Results and Analysis
-5.1 Main Results
-This section presents a comprehensive evaluation of CoR-Math-7B’s capabilities in diverse mathematical reasoning tasks.
-We systematically assess its performance against SOTA models and expert models on specific domains using a zero-shot evaluation setting.
-Our objective is to demonstrate how multi-paradigm approach enables CoR-Math-7B to achieve superior general mathematical reasoning abilities and computational efficiency, compared to methods based on a single reasoning paradigm.
-5.1.1 Comparisons with General-purpose Mathematical Models
-Table 1: A overall comparison of CoR-Math-7B with three types of general mathematical reasoners (proprietary, foundational, and general-purpose mathematical models (GMM)) on three mathematical benchmarks.
+Further details are in Sec.˜B.4.
+[…]
+Table 1: A overall comparison of CoR-Math-7B with five types of general mathematical reasoners (Proprietary, Foundational, and General-purpose Mathematical Models (GMM)) on three mathematical benchmarks.
 Results are shown for zero-shot (denoted by $\emptyset$) or few-shot settings by default.
@@
 Bolded scores are the best performance among all models except for the proprietary one.
-Model Model Types MATH GSM8k miniF2F-test Sample Budget ($N$) miniF2F-test Pass@1 Pass@1 Pass@$N$ o1-mini Proprietary 90.0 94.8 1 13.2 GPT-4 Proprietary 42.5 87.1 128 24.6 GPT-4o Proprietary 76.6 90.5 128 25.0 Llama-3.1-8B Foundation 4.2 6.2 128 25.8 Llama-3.1-8B-Instruct Foundation 47.2 76.6 128 23.4 Mistral-7B Foundation 14.3 40.3 1$\times$32$\times$100 22.1 Mixtral-8x7B Foundation 28.4 74.4 1$\times$32$\times$100 23.4 MUSTARD GMM 13.8 27.9 1 7.8 Llemma-7B GMM 18.6 41.0 1$\times$32$\times$100 26.2 DeepSeek-Math-7B-Base GMM 11.8 22.2 1$\times$32$\times$100 28.3 InternLM2-Math-7B-Base GMM 21.5 49.2 1$\times$32$\times$100 30.3 InternLM2-Math-Plus-7B GMM 53.0 85.8 1$\times$32$\times$100 43.3 CoR-Math-7B GMM 66.7 88.7 128$\times$1 52.9 32$\times$100 59.4 128$\times$128 66.0
-To explore the versatility of CoR-Math-7B and assess its comprehensive mathematical reasoning capabilities, we compare it with widely-used models and SOTA general mathematical models.
-As shown in Table 1, CoR-Math-7B achieves the best performance across three challenging benchmarks in a zero-shot setting.
-These results collectively highlight the strong comprehensive mathematical reasoning abilities of our model.
-The findings are detailed below: (1) arithmetic computation task: our CoR-Math-7B achieves optimal results across all arithmetic computation benchmarks.
-For example, it exceeds InterLM2-Math-Plus-7B by a $13.7\%$ absolute improvement on the MATH dataset.
-(2) theorem proving task: our CoR-Math-7B achieves the best zero-shot result on MiniF2F, even surpassing the proprietary GPT-4o model by a $41\%$ absolute increase in few-shot setting.
-(3) zero-shot vs few-shot: The zero-shot results of CoR-Math-7B outperform all few-shot results from other models.
-For instance, the zero-shot performance of CoR-Math-7B demonstrates a $14.0\%$ absolute improvement over the few-shot performance of DeepSeek-math-7B-Base on the MiniF2F dataset.
-When testing on the MiniF2f dataset revealed limited proficiency in Qwen2.5-Math (Yang et al., 2024) in Lean 4, indicating a focus on arithmetic computation over theorem proving.
-Therefore, we classify Qwen2.5-Math as an arithmetic computation model rather than a general mathematical model and compare it in Section 5.1.3.
-These findings indicate the strong generalization capability and comprehensive mathematical reasoning ability of the CoR framework.
+Metrics.
+Accuracy is the primary evaluation metric.
+For arithmetic tasks, we adopt widely-used CoT settings (Wei et al., 2022), rounding answers to the nearest integer.
+The SymPy library [^1] https://www.sympy.org/ is used for parsing and evaluation.
+To handle numerical representation variations, the model explicitly states the final answer (LI et al., 2024).
+For theorem proving, we follow the recent advances (Xin et al., 2024), adapting the miniF2F benchmark from Lean 3 to Lean 4.
+The pass@$N$ metric evaluates proof correctness within $N$ attempts.
+Our CoR-Math-7B uses SMPS with NLR and SR, where $N=N_{\text{NLR}}\times N_{\text{SR}}$.
+Model settings are detailed in Sec.˜B.5.
+4.2 Implementation Details
+We fine-tuned widely-used DeepSeekMath-Base 7B (Shao et al., 2024) and Llama-3.1 8B (Dubey et al., 2024) models, employing our PPT method on MPM dataset.
+Unless otherwise specified, CoR-Math-7B model is based on DeepSeekMath-Base 7B.
+The details are available in Sec.˜B.3.
+4.3 Baselines
+We examine three categories of baseline models, with results reported using CoT prompting.
+General-purpose mathematical models.
+To evaluate CoR’s generalization, we includes Mustard (Huang et al., 2024), DeepSeekMath (Shao et al., 2024), InternLM-Math (Ying et al., 2024b), Llama-3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Llemma (Azerbayev et al., 2024).
+Task-specific mathematical models.
+We consider several expert models on mathematical optimization for specific tasks, such as large-scale mathematical data and inference search improvements.
+The arithmetic experts encompass Qwen2.5-Math (Yang et al., 2024), WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2024), DART-Math (Tong et al., 2024), InternLM-Math (Ying et al., 2024b), , DeepSeekMath-Instruct / RL (Shao et al., 2024), Xwin-Math (Li et al., 2024), ToRA (Gou et al., 2024), and NuminaMath (LI et al., 2024).
+The theorem proving experts include LLM-Step (Welleck and Saha, 2023), GPT-f (Polu and Sutskever, 2020), Lean-STaR (Lin et al., 2024a), Hypertree Proof Search (Lample et al., 2022), DeepSeek-Prover (Xin et al., 2024), and InternLM2.5-StepProver (Wu et al., 2024).
+Foundation and proprietary models.
+We present open-source foundation models Llama-3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Mixtral (Jiang et al., 2024), along with proprietary models o1-mini (OpenAI, 2024), GPT-4 (OpenAI, 2023), and GPT-4o (Hurst et al., 2024).
+5 Main Results
+5.1 Comparisons with General-purpose Mathematical Models
+To evaluate the comprehensive mathematical reasoning abilities of CoR-Math-7B, we compare it with widely-used models and SOTA mathematical models, which are based on single reasoning paradigms.
+As shown in  Table˜1, CoR-Math-7B outperforms others across five challenging benchmarks in a zero-shot setting, highlighting its strong cross-task versatility.
+The main findings are: (1) arithmetic computation: CoR-Math-7B achieves the best results, outperforming InterLM2-Math-Plus-7B by an absolute $13.7\%$ on MATH and $2.9\%$ on GSM8K.
+Compared to Llama-3.1-8B-Instruct, it achieves an increase in correct answers on AMC2023 ($34$ vs $16$) and AIME2024 ($12$ vs $5$).
+(2) theorem proving: CoR-Math-7B sets a new zero-shot benchmark on miniF2F, even surpassing GPT-4o model by a $41.0\%$ absolute increase in few-shot setting.
+(3) zero-shot vs few-shot: CoR-Math-7B’s zero-shot performance exceeds all few-shot results from other models, with an absolute $37.7\%$ improvement over DeepSeekMath-7B-Base on miniF2F.
+These findings indicate the generalization capability and comprehensive mathematical reasoning ability of the CoR framework.
 This suggests NLR descriptions and SR verification rehearse precise mathematical reasoning for AR.
-In summary, the results indicate that the CoR framework can handle a diverse range of complex reasoning challenges, and leveraging the synergy between different paradigms can enhance general reasoning capabilities.
-5.1.2 Comparisons with Expert Models on Theorem Proving
-Table 2 shows CoR-Math-7B’s exceptional zero-shot performance on the MiniF2F, achieving $66.0\%$ accuracy without demonstrations while competing methods require few-shot examples.
-Its zero-shot performance is particularly noteworthy when considering computational efficiency.
-Table 2: The zero-shot ($\emptyset$) performance of CoR-Math-7B and the few-shot performance of optimized models on the MiniF2F benchmark.
-These models incorporate optimizations related to theorem proving.
-The highest scores are in bold.
-Model Sample Budget $N$ MiniF2F LLMStep 1$\times$32$\times$100 27.9 GPT-f 64$\times$8$\times$512 36.6 Hypertree Proof Search 64$\times$5000 41.0 Lean-STaR 64$\times$1$\times$50 46.3 DeepSeek-Prover-V1.5-Base 6400 42.2 DeepSeek-Prover-V1.5-SFT + RMaxTS 4$\times$6400 56.3 DeepSeek-Prover-V1.5-SFT + RMaxTS 32$\times$6400 60.2 DeepSeek-Prover-V1.5-RL + RMaxTS 4$\times$6400 59.6 DeepSeek-Prover-V1.5-RL + RMaxTS 32$\times$6400 63.5 InternLM2.5-StepProver 64$\times$32$\times$100 54.5 InternLM2.5-StepProver-BF 1$\times$32$\times$600 47.3 InternLM2.5-StepProver-BF 256$\times$32$\times$600 59.4 InternLM2.5-StepProver-BF+CG 2$\times$32$\times$600 50.7 InternLM2.5-StepProver-BF+CG 256$\times$32$\times$600 65.9 CoR-Math-7B 128$\times$128 66.0
-For instance, under similar computational constraints, our CoR-Math-7B surpasses DeepSeek-Prover-V1.5-RL + RMaxTS by $6.4\%$ absolute points ($66.0\%$ vs $59.6\%$ with $128\times 128$ vs $4\times 6400$) and InternLM2.5-StepProver-BF+CG by $15.3\%$ absolute points ($66.0\%$ vs $50.7\%$ with $128\times 128$ vs $2\times 32\times 600$).
-Compared to significantly larger sample sizes, CoR-Math-7B retains its competitive edge.
-For example, zero-shot CoR-Math-7B outperforms few-shot InternLM2.5-StepProver-BF+CG by $0.1\%$ in absolute score, despite considering approximately $300$ times more potential solutions.
-These results highlight the effectiveness of multi-paradigm reasoning and demonstrate how CoR efficiently searches paradigm-based solution spaces.
-5.1.3 Comparisons with Expert Models on Arithmetic Computation
-Table 3 shows the significant competitiveness of our CoR-Math-7B when compared to expert models on arithmetic computation.
-Compared to methods like ToRA and NuminaMath that use code as a tool without full reasoning, our CoR-Math-7B with multi-paradigm reasoning outperforms them on all benchmarks.
-For instance, on the competition-level MATH dataset, CoR-Math-7B achieves an $11.4\%$ absolute improvement over the tool-integrated NuminaMath-7B-TIR.
-This reinforces the effectiveness of complete code-based reasoning compared to the fragmented interleaving of natural language and Python code.
-To further explore the relationship between resource utilization (training sample size) and performance, we fit a quadratic function to the current SOTA models and conduct a Pareto frontier analysis (Branke et al., 2008).
-Based on Table 3, we generate the fitting Pareto curve shown in Figure 1 (b), which is based on all reasoning methods.
-Our CoR-Math-7B achieves superior performance compared to the optimal performance of single-paradigm reasoning methods with similar data volume.
-This suggests the existence of a new optimal curve based on multi-paradigm reasoning methods, which outperforms the curve of single-paradigm approaches.
+Additionally, we provide qualitative analysis of error cases in  Sec.˜D.2.
+In summary, CoR effectively handles diverse reasoning challenges, with synergies across paradigms boosting its overall performance.
+[…]
+Table 2: The zero-shot ($\emptyset$) performance of CoR-Math-7B and the few-shot performance of theorem proving optimized models on the miniF2F benchmark. The highest scores are in bold.
+5.2 Comparisons with Theorem Proving Experts
+Table˜2 shows CoR-Math-7B’s impressive zero-shot performance on miniF2F, achieving $66.0\%$ accuracy without demonstrations, while competitors require few-shot examples.
+Its zero-shot performance is noteworthy when considering computational efficiency.
+For instance, under similar computational constraints, CoR-Math-7B surpasses DeepSeek-Prover-V1.5-RL + RMaxTS by $6.4\%$ absolute points ($66.0\%$ vs $59.6\%$ with $128\times 128$ vs $4\times 6400$) and InternLM2.5-StepProver-BF+CG by $15.3\%$ absolute points ($66.0\%$ vs $50.7\%$ with $128\times 128$ vs $2\times 32\times 600$).
+Even with larger sample sizes, CoR-Math-7B remains superior, surpassing few-shot InternLM2.5-StepProver-BF+CG by $0.1\%$, despite considering $300$ times more solutions.
+These results highlight the efficiency of multi-paradigm reasoning and CoR’s ability to explore paradigm-based solution spaces.
 […]
 […]
-5.2 Ablation Study
-We conduct a series of ablation studies, providing insights into the key factors that drive the performance of our multi-paradigm reasoning framework.
-5.2.1 Impact of Stages in PPT Method
-Figure 4: An evaluation of the effectiveness of the PPT strategy on mathematical reasoning tasks.
-We present the zero-shot Pass@1 results on the MATH and GSM8k benchmarks, across three cumulative stages of the PPT strategy based on the DeepSeek-Math and Llama-3.1 models.
-.
-The results highlight the cumulative effectiveness of the PPT strategy, showing increased performance for both models across all benchmark datasets with each progressive stage.
-These results indicate that the PPT strategy is a viable method to enhance mathematical reasoning capabilities.
-To evaluate the effectiveness of PPT methods, we employ two base models, Llama-3.1-8B (Dubey et al., 2024) and Deepseek-Math-7B-Base (Shao et al., 2024).
-We present the results of the PPT method across different stages on these two models for the MATH and GSM8K benchmarks.
-As shown in Figure 4, the vanilla models exhibit limited performance on both benchmarks.
-After stage ①, the performance of all models improves significantly.
-For instance, Llama-3.1-Base achieves absolute gains of $47.9\%$ on MATH and $61.0\%$ on GSM8K.
-This demonstrates the necessity of understanding mathematical concepts and warming up the models.
-Subsequently, we conduct stage ② training, which utilizes data with NLR and AR.
-The limited performance gains are potential because the pre-training data already includes code and natural language corpus.
-Both stage ① and stage ② primarily reactivate the base models in the mathematical domain, and stage ① already incorporates NLR training.
-In the final stage ③, we consider training with three paradigms, and all models experience performance enhancements.
-This suggests that introducing rare or unseen paradigms helps to improve the reasoning ability of learned paradigms.
-5.2.2 Impact of Reasoning Paradigms Order
-We investigated the impact of varying the sequence of reasoning paradigms employed by CoR-Math-7B during zero-shot inference for arithmetic computation tasks.
-Our experiments involved modifying only the prompt structure to enforce different paradigm orders.
-Given that NLR closely aligns with the pre-training of LLMs, it was consistently positioned as the initial paradigm in our experiments.
-Table 4 presents the accuracy for the MATH and GSM8K benchmarks under two different paradigm orders.
-The results indicate that performing SR before AR achieves the best performance.
-Specifically, on the MATH benchmark, this order achieves an accuracy of $66.7\%$, an improvement over the $49.9\%$ attained by the sequence that AR before SR.
-A possible explanation for this observation lies in the synergistic effect of the ordered paradigms.
-SR plays a crucial role in decomposing the problem into manageable sub-steps, thereby providing a structured foundation for subsequent AR.
-By placing SR before AR, we hypothesize that the model is better equipped to perform accurate calculations, leveraging the intermediate reasoning steps generated by SR.
-Furthermore, positioning AR immediately preceding the Summary phase may facilitate a more coherent integration of the computational outcome into the final answer.
+5.3 Comparisons with Arithmetic Experts
+Since several mathematical reasoners are primarily trained on arithmetic computation rather than formal theorem proving, and often struggle with the latter, we classify them as experts in arithmetic computation to enable a fair comparison.
+Table˜3 highlights CoR-Math-7B’s strong performance in arithmetic computation, surpassing expert models and achieving a $15\%$ improvement over RL-based methods, with DeepSeekMath-RL-7B as the representative, on the MATH benchmark for arithmetic tasks.
+Unlike methods such as ToRA and NuminaMath, which rely on code without complete reasoning, CoR-Math-7B leverages multi-paradigm reasoning to outperform them across all benchmarks.
+For example, on the MATH dataset, CoR-Math-7B achieves an $11.4\%$ absolute improvement over the tool-integrated NuminaMath-7B-TIR, underscoring the effectiveness of complete code-based reasoning over fragmented interleaving of natural language and code.
+To further explore the relationship between resource utilization (training sample size) and performance, we fit a quadratic function to the current SOTA models and perform a Pareto frontier analysis (Branke et al., 2008), as shown in Fig.˜1 (Details in Sec.˜B.6).
+CoR-Math-7B outperforms the optimal performance of single-paradigm methods with similar data volume, suggesting a new optimal curve for multi-paradigm reasoning that surpasses single-paradigm approaches.
+6 Ablation Study
+Figure 4: An evaluation of the effectiveness of the PPT strategy.
+We present the zero-shot Pass@1 results on the MATH and GSM8k benchmarks across three cumulative stages of the PPT strategy.
+The results highlight the PPT strategy’s cumulative effectiveness, showing increased performance with each progressive stage.
+6.1 Impact of Stages in PPT Method
+We evaluate the PPT method using two base models, Llama-3.1-8B and DeepseekMath-7B-Base, on the MATH and GSM8K benchmarks at different PPT stages.
+As shown in Fig.˜4, the vanilla models exhibit limited performance.
+After stage ①, performance improves significantly, with Llama-3.1-Base gaining $47.9\%$ on MATH and $61.0\%$ on GSM8K, emphasizing the need for mathematical understanding and model warming.
+Stage ②, utilizing NLR and AR data, yields limited gains compared to stage ① due to pre-training data already containing relevant content.
+However, this does not imply that stage ② is unimportant, as it plays a critical role in CoR by reactivating and refining computational skills essential for collaborating with other reasoning paradigms.
+Both stages ① and ② focus on reactivating models for mathematical tasks.
+In stage ③, training with three paradigms further enhances performance, suggesting that rare or unseen paradigms improve reasoning abilities.
+The consistent performance gains with added paradigms in  Fig.˜4 robustly demonstrate the superiority of our multi-paradigm PPT approach over single- or fewer-paradigm fine-tuning of the same base model, yielding a more powerful, comprehensive mathematical reasoner.
+Benchmark
+[…]
+SR $\rightarrow$ Summary
+[…]
+AR $\rightarrow$ Summary
+MATH 49.9 66.7
+GSM8K 84.2 88.7
 Table 4: Zero-shot Pass@1 results for varying paradigm orders on the MATH and GSM8k benchmarks. The best results are in bold.
-[…]
-MATH 66.7 49.9
-GSM8K 88.7 84.2
-5.2.3 Impact of Scaling Model Size
-Table 5: Zero-shot Pass@1 results on the MATH and GSM8k benchmarks across different model scales. The best scores are in bold.
-Model MATH GSM8k
-Qwen2.5-Math-1.5B (CoR) 57.6 84.5
-Qwen2.5-Math-7B (CoR) 64.7 90.0
-Table 5 investigates the performance of various base models with different parameters under the CoR framework.
-We employ the Qwen2.5-Math models (Yang et al., 2024) with $1.5$ billion and $7$ billion parameters as the base model.
-Results demonstrate that our CoR framework achieves superior performance with increasing parameter scale.
-Specifically, when comparing a model with $7$ billion parameters to one with $1.5$ billion parameters, the former attained an absolute increase of $7.1\%$ on the MATH benchmark and $5.5\%$ on the GSM8K benchmark.
-This evidence supports the claim that CoR exhibits scalability with respect to model parameter scaling.
-6 Discussion on Reasoning Hierarchy: Paradigms, Paths, and Steps
+Model GSM8k@1 MATH@1 AMC2023 (Maj@64) AIME2024 (Maj@64) miniF2F@128 DSM + NLR 33.9 27.8 14/40 1/30 - DSM + AR 75.8 37.6 10/40 0/30 - DSM + SR - - - - 44.3 DSM + CoR (ours) 88.7 66.7 34/40 12/30 52.9
+Table 5: Zero-shot results of multi- and single-paradigm fine-tuning on DeepSeekMath-7B-Base (DSM) for mathematical Reasoning benchmarks. The best results are in bold.
+Model Size MATH GSM8k miniF2F
+Qwen2.5-Math (Base) 1.5B 34.0 39.3 0.0
+7B 51.8 90.0 0.0
+Qwen2.5-Math (CoR) 1.5B 57.6 84.5 51.6
+7B 64.7 90.0 52.5
+Llama-3.1 (Base) 8B 4.2 6.2 25.8
+70B 16.8 20.5 22.5
+Llama-3.1 (CoR) 8B 58.2 84.0 53.3
+70B 70.7 90.0 56.2
+Table 6: Zero-shot Pass@1 results on the mathematical benchmarks across different model scales. The best scores are in bold.
+6.2 Order of the Reasoning Paradigms
+As shown in Table˜4, we investigate the impact of varying the sequence of reasoning paradigms in CoR-Math-7B during zero-shot arithmetic tasks, modifying only the prompt structure.
+Given that NLR closely aligns with the pre-training of LLMs, it serves as the first paradigm.
+The results indicate that SR before AR resulted in the highest accuracy, with $66.7\%$ on MATH, compared to $49.9\%$ for AR before SR.
+A possible explanation is that SR plays a crucial role in decomposing the problem into manageable sub-steps, thereby providing a structured foundation for subsequent AR.
+This demonstrates that paradigms in CoR are interdependent, not isolated, and performance hinges on how they are linked during reasoning, beyond merely mastering individual paradigms.
+Furthermore, positioning AR preceding the Summary boxes may facilitate a more coherent integration of the computational outcome into the final answer.
+6.3 Impact of Model Scales
+Table˜6 evaluates base model performance under the CoR framework with varying parameters, such as Qwen2.5-Math models (1.51.5B / 77B) and Llama-3.1-8B (88B / 7070B).
+The results show that CoR scales with model size.
+The 77B Qwen2.5-Math outperforms the 1.51.5B version by $7.1\%$ on MATH and $5.5\%$ on GSM8K.
+Similarly, Llama-3.1 7070B gains $2.9\%$ over its 88B version on minF2F.
+This supports that CoR exhibits parameter scalability.
+6.4 Multi- vs. Single-Paradigm Reasoning
+To further validate the superior performance of multi-paradigm reasoning, we conducted an ablation experiment comparing it to single-paradigm fine-tuning We fine-tuned the DeepSeekMath-7B-Base model separately using single-paradigm paths (NLR, AR, SR) extracted from the MPM dataset.
+All experiments employed identical base models, problem sets, and training data sizes ($10,000$ samples per paradigm).
+Experimental results, summarized in Table 5, clearly demonstrate that our CoR model significantly outperforms single-paradigm fine-tuned models across diverse tasks.
+Specifically, for GSM8k and MATH datasets, CoR achieves improvements of $12.9\%$ and $29.1\%$, respectively, compared to the best-performing single-paradigm approaches.
+On AMC2023 and AIME2024, CoR solves $20$ and $11$ more problems, respectively, compared to single-paradigm approaches.
+Lastly, on minF2F, CoR achieves an accuracy improvement of $8.6\%$.
+Collectively, these improvements underline CoR’s effectiveness in integrating multiple reasoning paradigms, enabling robust and generalized problem-solving capabilities.
+7 Conclusions
+This paper introduces CoR, a novel unified reasoning framework that enhances mathematical reasoning in LLMs by synergistically integrating natural language, algorithmic, and symbolic.
+CoR tackles the limitations of single-paradigm approaches with the PPT strategy and the MPM dataset, enabling LLMs to progressively master diverse reasoning paradigms for improved generalization and performance.
+Our CoR-Math-7B outperformed SOTA models on five challenging benchmarks, showcasing enhanced zero-shot generalization.
+Ablation studies validated the benefits of progressive training and paradigm sequence.
+Additionally, CoR offers a new perspective on test-time scaling through multi-paradigm reasoning to enhance both efficiency and performance.
+Acknowledgments
+This work was partly supported by the Shenzhen Science and Technology Program ( JSGG20220831110203007 ) and the research grant No. CT20240905126002 of the Doubao Large Model Fund.
+Limitations
+This paper proposes a general zero-shot reasoning method that supports different paradigms to solve multiple tasks.
+Evaluating zero-shot performance is challenging, as many methods rely on a single evaluation metric, like zero-shot pass@1.
+Additionally, most approaches focus on few-shot settings, as seen in benchmarks like miniF2F, limiting the comprehensiveness of our evaluation with aligned settings.
+To address this, we include additional experiments in Sec.˜C.2 to examine the impact of different evaluation strategies.
+On the other term, CoR can explore a multi-paradigm solution space, covering up to three paradigms, such as in large-scale arithmetic tasks.
+However, due to resource constraints and the fact that other methods can only predict within a single paradigm, we did not further consider using SMPS on benchmarks such as MATH to ensure fairness.
+Appendix
+Appendix A Discussion on Reasoning Hierarchy: Paradigms, Paths, and Steps
 This section attempts to explore the essence of reasoning, contrasts CoR with current methods, examines the factors contributing to its effectiveness, and provides a new perspective for future research.
-Figure 5: An overview of the reasoning hierarchical structure.
-(a) A single reasoning paradigm depicts multiple distinct reasoning paths.
-(b) An example of single-paradigm reasoning includes one reasoning path, which contains several reasoning steps.
-(c) Multi-paradigm reasoning includes several distinct reasoning paradigms.
-As shown in Figure 5, this paper posits that reasoning texts generated by LLMs exhibit a reasoning hierarchical structure, which consists of three levels: reasoning paradigms, reasoning paths, and reasoning steps.
+As shown in Fig.˜5, this paper posits that reasoning texts generated by LLMs exhibit a reasoning hierarchical structure, which consists of three levels: reasoning paradigms, reasoning paths, and reasoning steps.
   • Reasoning steps represent the fundamental units, each comprising one or more tokens and encompassing an incomplete stage of the solution process.
@@
 They often contain multiple potential final answers, thus necessitating a summarization method, such as a summary module, to extract the ultimate answer.
-Furthermore, a reasoning paradigm utilizes a single knowledge media, such as natural language.
-Figure 6: A comparison of reasoning paradigms in advanced studies.
-(a) deep reasoning, characterized by serial concatenation of reasoning paths.
-(b) interleaved reasoning, which integrates secondary methods for guidance within a dominant paradigm.
-(c) multi-paradigm reasoning (Our proposed CoR), a framework designed to harness the synergistic effects between different reasoning paradigms.
+Furthermore, a reasoning paradigm uses a single knowledge media, such as natural language.
+Figure 5: An overview of the reasoning hierarchical structure.
+(a) A single reasoning paradigm depicts multiple distinct reasoning paths.
+(b) An example of single-paradigm reasoning includes one reasoning path, which contains some reasoning steps.
+(c) Multi-paradigm reasoning includes several distinct reasoning paradigms.
+Figure 6: A comparison of reasoning methods in several advanced studies.
 Furthermore, current studies can be categorized based on their focus.
-As shown in Figure 5 (a), contemporary work concentrates within a single paradigm, optimizing along two dimensions: depth (the number of reasoning steps) and width (the number of reasoning paths).
+As shown in Fig.˜5 (a), contemporary work concentrates within a single paradigm, optimizing along two dimensions: depth (the number of reasoning steps) and width (the number of reasoning paths).
 For instance, regarding reasoning depth, the CoT method (Wei et al., 2022) employs prompts to increase intermediate steps within one reasoning path to achieve higher performance.
@@
 Building upon these generated reasoning paths, existing work proposes various integration strategies to obtain the final answer, such as Best-of-N and Self-consistency (Wang et al., 2023).
-Specifically, as shown in Figure 6 (a), recent advanced studies introduce the deep reasoning method, which focuses on generating serial concatenation of reasoning paths followed by summarization (like OpenAI o1).
+Specifically, as shown in Fig.˜6 (a), recent advanced studies introduce the deep reasoning method, which focuses on generating serial concatenation of reasoning paths followed by summarization (like OpenAI o1 OpenAI (2024)).
 Since these methods are based on different optimization dimensions, we can combine them feasibly in applications.
 Specifically, some approaches prioritize a dominant paradigm for reasoning yet hope to integrate other paradigms for guidance.
-For instance, as shown in Figure 6 (b), the reasoning process of ToRA (Gou et al., 2024) presents an interleaved reasoning method with different knowledge media.
+For instance, as shown in Fig.˜6 (b), the reasoning process of ToRA (Gou et al., 2024) presents an interleaved reasoning method with different knowledge media.
 It involves initial natural language generation, which is followed by code generation.
@@
 Current methods often overlook the synergistic effects between paradigms and underestimate the importance of complete reasoning for achieving accurate results.
-To address this challenge, as shown in Figure 5 (c) and Figure 6 (c), we introduce Chain-of-Reasoning (CoR), a unified reasoning framework capable of multi-paradigm reasoning.
+To address this challenge, as shown in Fig.˜5 (c) and Fig.˜6 (c), we introduce Chain-of-Reasoning (CoR), a unified reasoning framework capable of multi-paradigm reasoning.
 This framework embodies the following potential advantages:
@@
 For instance, knowledge derived from a natural language paradigm guides subsequent algorithmic paradigms, thereby enhancing the efficiency and accuracy of code generation based on detailed natural language reasoning.
-This mechanism also provides substantial context from prior paradigms to later paradigms and operates as a form of preliminary rehearsal where subsequent paradigms can follow correct reasoning steps or rectify incorrect ones.
+This mechanism also provides substantial context from prior paradigms to later paradigms.
+It operates as a form of preliminary rehearsal where subsequent paradigms can follow correct reasoning steps or rectify incorrect ones.
 A novel direction for test-time scaling emerges.
@@
 This implies that CoR serves as a platform that facilitates continuous integration and collaboration.
-7 Conclusions
-This paper introduces Chain-of-Reasoning (CoR), a novel unified reasoning framework that synergistically combines natural language, algorithmic, and symbolic reasoning to solve mathematical problems.
-Our approach demonstrates the potential of integrating diverse reasoning paradigms in the reasoning process, enabling language models to tackle complex mathematical challenges effectively.
-CoR achieves significant performance gains on challenging mathematical reasoning tasks, substantially outperforming existing single-paradigm approaches on benchmarks like MATH and MiniF2F.
-In addition, CoR introduces a new, orthogonal perspective to traditional test-time scaling by leveraging multi-paradigm reasoning paths to enhance computational efficiency and performance.
-Furthermore, our proposed CoR framework offers valuable insights for future research, contributing to the development of more robust and versatile reasoning agents capable of leveraging the strengths of multiple reasoning modalities.
-Appendix
-Appendix A Experiment Details
-A.1 The detail of the Universal Text Template
+Appendix B Experiment Details
+In this study, we use open-source tools, models, and datasets in compliance with their respective open-source licenses and solely for academic research purposes.
+B.1 The detail of the Universal Text Template
 Universal Template Designed for Multi-paradigm Reasoning.
@@
 Figure 7: The universal text template designed for multi-paradigm reasoning. The contents for each paradigm have been omitted.
-As shown in Figure 7, we apply the designed universal text template on all training samples.
-A.2 Example Prompts for Dataset Enhancement
+As shown in Fig.˜7, we apply the designed universal text template on all training samples.
+B.2 Example Prompts for Dataset Enhancement
 To guide LLMs in generating and refining reasoning paradigms during dataset enhancement, we developed specific prompts ($p_{s}$) tailored to each seed dataset.
 These prompts provide structured instructions for the models to follow, ensuring the generation of high-quality and logically consistent reasoning paradigms.
-Figure 9 and Figure 8 provide examples of such prompts designed for augmenting and refining formal proofs within the Lean 4 theorem prover for the Numina-TIR and Lean-Workbook datasets, respectively.
+Fig.˜8 and Fig.˜9 provide examples of such prompts designed for augmenting and refining formal proofs within the Lean 4 theorem prover for the Numina-TIR and Lean-Workbook datasets, respectively.
 The prompt specifies the input and output format and includes placeholders for the problem statement, informal proof, and corresponding formal proof in Lean 4.
 It also incorporates $N_{\text{shot}}$ few-shot examples to further guide the model.
-In our experiments, consistent with the common settings employed in (Jiang et al., 2023; Yang et al., 2024), we use $N_{\text{shot}}=8$.
-This structured approach and setting are adapted and applied to all datasets to ensure comprehensive coverage and logical consistency across the enhanced dataset.
+In our experiments, we follow the common settings (Jiang et al., 2023; Yang et al., 2024; Wang et al., 2025; Zhao and Zhang, 2024) and set $N_{\text{shot}}=8$.
+This structured approach and setting are applied to all datasets to ensure comprehensive coverage and logical consistency across the enhanced dataset.
+B.3 Details of Training Settings
+In all stages of the PPT method, a learning rate of 2e52e-5 and a warm-up ratio of $1\%$ are implemented.
+To enhance computational efficiency, the training process is conducted using distributed optimization with DeepSpeed ZeRO Stage 3 (Rajbhandari et al., 2020) and is combined with Flash-Attention (Dao, 2024).
+Stage ➀ comprises $3$ epochs, whereas the stage ➁ and ➂ are conducted over $4$ epochs.
+A sequence length of $2,048$ tokens is used in Stages ➀ and ➁, which is increased to $4,096$ tokens in stage ➂ to support more complex reasoning tasks.
+Furthermore, we employ an annealing strategy at the end of stage ➂ with high-quality MPM samples, with the aim of further enhancing model accuracy in complex reasoning tasks.
+B.4 Benchmarks
+We report the statistics of the evaluation datasets in Table˜7.
+Dataset # of Test set Avg. Length
+MATH 5,000 30.7
+GSM8K 1,319 46.3
+AIME2024 30 59.2
+AMC2023 40 47.2
+MiniF2f 244 30.5
+Table 7: The statistics of the evaluation datasets.
+B.5 Details of Metrics
+This study employs the pass@$N$ metric on the miniF2F benchmark, with the evaluation grounded in a sample budget of $N$.
+To ensure a fair comparison of computational cost across different generation schemes, this paper defines the sample budget according to the following rules.
+  • For single-pass sampling methods, we define the sample budget $N$ as the total number of proofs generated, with large values of $N$ factorized for ease of comparison to tree search methods.
+  • For best-first-search methods, following the notation in Llemma (Azerbayev et al., 2024), we present $N=N_{\text{att}}\times N_{\text{tact}}\times N_{\text{iter}}$ where $N_{\text{att}}$ denotes the number of best-first-search attempts, $N_{\text{tact}}$ denotes the number of tactics generated for each expansion, and $N_{\text{iter}}$ denotes the number of expansion iterations.
+  • For tree-based search methods, e.g., RMaxTS (Xin et al., 2024) and HTPS (Lample et al., 2022), we present $N=N_{\text{att}}\times N_{\text{ex}}$, where $N_{\text{att}}$ denotes the number of tree search attempts, and $N_{\text{ex}}$ denotes the number of model generations invoked in tree expansions.
+  • For our SMPS, we define the sample budget as $N=N_{\text{NLR}}\times N_{\text{SR}}$, where $N_{\mathrm{NLR}}$ denotes the number of initial semantic reasoning paths, and $N_{\mathrm{SR}}$ denotes the number of symbolic reasoning paths extended for each other reasoning paradigm.
+[…]
+[…]
 Example Prompts for TIR Dataset Enhancement System Prompt:
@@
 Figure 9: Example pompts for Lean-Workbook dataset enhancement. The few-shot examples provided for the LLM have been omitted.
-A.3 Details of Training Settings
-In all stages of the PPT method, a learning rate of $52e-$ 5 and a warm-up ratio of $1\%$ are implemented.
-To enhance computational efficiency, the training process is conducted using distributed optimization with DeepSpeed ZeRO Stage 3 (Rajbhandari et al., 2020) and is combined with Flash-Attention (Dao, 2024).
-Stage ➀ comprises $3$ epochs, whereas the satge ➁ and ➂ are conducted over $4$ epochs.
-A sequence length of $2,048$ tokens is used in Stages ➀ and ➁, which is increased to $4,096$ tokens in Stage ➂ to support more complex reasoning tasks.
-Furthermore, we employ an annealing strategy at the end of Stage ➂ with high-quality MPM samples, with the aim of further enhancing model accuracy in complex reasoning tasks.
-A.4 Benchmarks
-We report the statistics of the evaluation datasets in Table 6.
-Table 6: The statistics of the evaluation datasets.
-Dataset # of Test set Avg. Length
-MATH 5,000 30.7
-GSM8K 1,319 46.3
-MiniF2f 244 30.5
-A.5 Details of Metrics
-This study employs the pass@$N$ metric on the MiniF2F benchmark, with the evaluation grounded in a sample budget of $N$.
-To ensure a fair comparison of computational cost across different generation schemes, this paper defines the sample budget according to the following rules.
-  • For single-pass sampling methods, we define the sample budget $N$ as the total number of proofs generated, with large values of $N$ factorized for ease of comparison to tree search methods.
-  • For best-first-search methods, following the notation in Llemma (Azerbayev et al., 2024), we present $N=N_{\text{att}}\times N_{\text{tact}}\times N_{\text{iter}}$ where $N_{\text{att}}$ denotes the number of best-first-search attempts, $N_{\text{tact}}$ denotes the number of tactics generated for each expansion, and $N_{\text{iter}}$ denotes the number of expansion iterations.
-  • For tree-based search methods, e.g., RMaxTS (Xin et al., 2024) and HTPS (Lample et al., 2022), we present $N=N_{\text{att}}\times N_{\text{ex}}$, where $N_{\text{att}}$ denotes the number of tree search attempts, and $N_{\text{ex}}$ denotes the number of model generations invoked in tree expansions.
-  • For our SMPS, we define the sample budget as $N=N_{\text{NLR}}\times N_{\text{SR}}$, where $N_{\mathrm{NLR}}$ denotes the number of initial semantic reasoning paths, and $N_{\mathrm{SR}}$ denotes the number of symbolic reasoning paths extended for each other reasoning paradigm.
-Appendix B Additional Analysis
-B.1 The Risk of Data Leakage
+B.6 Experimental Details of MATH and GSM8K benchmarks
+For arithmetic-related results, we analyze the number of mathematical samples utilized by the baseline model following the pre-training phase, as illustrated in Fig.˜1.
+Specifically, as shown in Table˜8, we present the sample quantity along with the model performance.
+Appendix C Additional Analysis
+C.1 The Risk of Data Leakage
 We measure the Levenshtein distance (Miller et al., 2009) between problems in the training dataset and those in the test dataset to mitigate the risk of data leakage.
@@
 As a result, approximately $1,000$ cases, accounting for less than $0.6\%$ of our training dataset, were removed to prevent potential leakage.
-B.2 Different Evaluation Strategies on Arithmetic Benchmarks
+Model MATH GSM8k Pass@1 Maj@8 Pass@1 Maj@8 Llama-3.1-8B 58.18 59.46 84.00 87.26 DeepSeekMath-Base-7B 66.74 71.70 88.70 91.43
 Table 9: Zero-shot results of different evaluation strategies across MATH and GSM8k benchmarks.
-Model MATH GSM8k
-Pass@1 Maj@8 Pass@1 Maj@8
-Llama-3.1-8B 58.18 59.46 84.00 87.26
-DeepSeek-Math-Base-7B 66.74 71.70 88.70 91.43
-Table 7 explores the impact of various evaluation strategies on the CoR framework.
+C.2 Different Evaluation Strategies on Arithmetic Benchmarks
+Table˜9 explores the impact of various evaluation strategies on the CoR framework.
 The majority vote strategy benefits the CoR framework.
 Specifically, a majority vote strategy with $8$ candidate samples on GSM8K can exceed GPT-4o’s Pass@1 performance ($90.5\%$).
+C.3 Ethical Considerations
+In this work, we present a method that achieves robust zero-shot multi-paradigm reasoning capabilities.
+However, this comes with potential risks, which include two primary social and ethical concerns: 1) the risk of tool misuse, and 2) pre-existing risks embedded in the backbone model parameters.
+The CoR framework supports various reasoning paradigms, some of which are deeply integrated with specific tools.
+For instance, the AR paradigm may involve unvetted Python code libraries.
+Therefore, we recommend performing pre-deployment audits on tools that may be utilized by the model.
+Furthermore, our method is compatible with existing LLMs, and several studies have highlighted the presence of biased behaviors in models like LLaMa (Kong et al., 2024).
+To mitigate ethical risks, we encourage the use of risk-free language models during deployment, which can reduce potential harm.
+Recent research suggests incorporating ethical alignment processes, such as filtering training data or generated content, to minimize unnecessary risks Yu et al.
+(2023).
+Overall, we advocate for open discussions regarding the use of our framework, as a transparent and public environment can reduce the likelihood of misuse.
+Appendix D Case Studies
+D.1 Cases of Instruction-free Reasoning
+In the absence of explicit instructions, as illustrated in Fig.˜10, the model generally adopts a paradigm sequence that aligns with the task types present in the training data, as described by MPM.
+The model transitions between paradigms autonomously.
+For example, during a computational task, the model initiates reasoning with the paradigm NLR, proceeds to SR for logical verification, and ultimately employs AR to achieve enhanced computational performance.
+D.2 Qualitative Analysis of Error Cases
+We now present a preliminary error pattern analysis for CoR-Math-7B across two tasks: arithmetic computation and theorem proving.
+In arithmetic computation, an analysis of 200 randomly sampled error cases reveals four primary types of errors: comprehension errors or incorrect premise assumptions (33 cases), logical fallacies (26 cases), computational errors (56 cases), and incomplete proofs or code syntax errors, which were most prevalent in the AR/SR paradigms (85 cases).
+Notably, cross-paradigm self-correction behaviors were observed.
+For example, AR rectified erroneous calculations initially produced by NLR, and SR supplemented incomplete logical steps that were missing in NLR outputs.
+In theorem proving, a separate analysis of 50 randomly sampled error cases revealed the same four error types, with the following distribution: 2 comprehension errors, 6 logical fallacies, 9 computational errors, and 33 incomplete proofs or syntax errors.
+Cross-paradigm self-correction behaviors were again consistently observed.
+As shown in  Fig.˜11 and  Fig.˜12, we provide an example of error cases and an example of self-correction behavior.
+We attribute these errors to the following potential causes.
+Despite MPM’s systematic integration of multi-paradigm reasoning paths and PPT’s paradigm-progressive training as discussed in  Sec.˜3.2 and  Sec.˜3.4 to mitigate distribution shifts, the overwhelming prevalence of NLR examples in the pretraining data —particularly the scarcity of SR-specific instances —has resulted in persistent competency gaps across paradigms.
+Additionally, CoR inherits the base LLM’s fixed-context window, which introduces the risk of truncating long reasoning chains or omitting key steps, leading to incomplete outputs.
+Furthermore, the observed self-correction capabilities can be attributed to the complementarity of reasoning paradigms within CoR, where one reasoning paradigm can rectify errors made by another, thus validating the effectiveness of our multi-paradigm approach.
+Additionally, feedback from multi-step reasoning validation allows subsequent paradigms to verify earlier outputs, facilitating error detection and correction throughout the reasoning process.
+Examples of Instruction-free Reasoning Generated by CoR Problem:
+Cities A and B are 45 miles apart.
+Alicia lives in A and Beth lives in B.
+Alicia bikes towards B at 18 miles per hour.
+Leaving at the same time, Beth bikes toward A at 12 miles per hour.
+How many miles from City A will they be when they meet?
+Solution:
+Let’s go through this step-by-step in natural language.
+To determine how many miles from City A Alicia and Beth will meet, we can use the concept of relative speed and the distance between their homes.
+1.
+Calculate the relative speed between Alicia and Beth: - Alicia’s speed: 18 miles per hour - Beth’s speed: 12 miles per hour - Relative speed = 18 + 12 miles per hour 2.
+Calculate the time it takes for them to meet: - Distance between City A and City B: 45 miles - $\text{Time to meet}=\text{Distance}\div\text{Relative speed}$ 3.
+Calculate the distance each person covers before meeting: - $\text{Distance covered by Alicia}=\text{Speed of Alicia}\times\text{Time to meet}$ - $\text{Distance covered by Beth}=\text{Speed of Beth}\times\text{Time to meet}$.
+Next, let’s write the corresponding formal proof in Lean 4 to prove this.
+# Formal proof in Lean 4:
+‘‘‘lean4
+def distance_AB := 45  -- miles
+def speed_Alicia := 18  -- miles per hour
+def speed_Beth := 12  -- miles per hour
+-- Calculate the relative speed
+def relative_speed := speed_Alicia + speed_Beth  -- miles per hour
+-- Calculate the time to meet
+def time_to_meet := distance_AB.toFloat / relative_speed.toFloat  -- in hours
+-- Calculate the distance each person covers before meeting
+def distance_covered_by_Alicia := speed_Alicia.toFloat * time_to_meet
+def distance_covered_by_Beth := speed_Beth.toFloat * time_to_meet
+-- Calculate how many miles from City A they will be when they meet
+def distance_from_A_when_meet := distance_covered_by_Alicia
+#eval distance_from_A_when_meet  -- returns 27.0
+‘‘‘lean_output
+27.000000
+Let’s use Python code to perform these calculations. # Code in Python:
+‘‘‘python
+import sympy as sp
+# Given values
+distance_AB = 45  # miles
+speed_Alicia = 18  # miles per hour
+speed_Beth = 12  # miles per hour
+# Calculate the relative speed
+relative_speed = speed_Alicia + speed_Beth  # miles per hour
+# Calculate the time to meet
+time_to_meet = distance_AB / relative_speed  # in hours
+# Calculate the distance each person covers before meeting
+distance_covered_by_Alicia = speed_Alicia * time_to_meet
+distance_covered_by_Beth = speed_Beth * time_to_meet
+# Calculate how many miles from City A they will be when they meet
+distance_from_A_when_meet = distance_covered_by_Alicia
+print(distance_from_A_when_meet)
+‘‘‘output
+27.0
+Alicia and Beth will meet boxed27 miles from City A.
+Figure 10: A case of the reasoning trajectory produced by CoR in an instruction-free setting.
+Examples of Incomplete Proofs or Code Syntax Errors Problem:
+What is the value of $\sqrt{1{,}\!000{,}\!000}-\sqrt[3]{1{,}\!000{,}\!000}$? Show that it is 900.
+Solution:
+Let’s go through this step-by-step in natural language.
+We need to prove the following theorem:
+theorem mathd_algebra_208 : Real.sqrt 1000000 - 1000000 ˆ ((1 : ℝ) / 3) = 900 := by
+We have the following: $\sqrt{1,000,000}-\sqrt[3]{1,000,000}$ $=\sqrt{10^{6}}-\sqrt[3]{10^{6}}$ $=(10^{6})^{\frac{1}{2}}-(10^{6})^{\frac{1}{3}}$ $=10^{6\cdot\frac{1}{2}}-10^{6\cdot\frac{1}{3}}$ $=10^{3}-10^{2}=1000-100=900.$ So, the value of $\sqrt{1,000,000}-\sqrt[3]{1,000,000}$ is indeed 900.
+Next, let’s write the corresponding formal proof in Lean 4 to prove this.
+# Formal proof in Lean 4:
+‘‘‘lean4
+theorem mathd_algebra_208 : Real.sqrt 1000000 - 1000000 ˆ ((1 : ℝ) / 3) = 900 := by
+norm_num [sqrt_eq_iff_mul_self_eq_of_pos, pow_one]
+‘‘‘lean_output
+Execution Failed.
+The original code has a syntax error. The proof block after := by is incomplete. Proof failed.
+Figure 11: A proof task failure case caused by incomplete proofs or code syntax errors.
+Examples of Instruction-free Reasoning Generated by CoR Problem:
+Let $b\geq 2$ be an integer. Call a positive integer $n$ $b$-eautiful if it has exactly two digits when expressed in base $b$, and these two digits sum to $\sqrt{n}$.
+For example, $81$ is $13$-eautiful because $81=\underline{6}\,\underline{3}_{13}$ and $6+3=\sqrt{81}$.
+Find the least integer $b\geq 2$ for which there are more than ten $b$-eautiful integers.
+Solution:
+Let’s go through this step-by-step in natural language. To solve this problem, we need to find the smallest integer $b\geq 2$ such that there are more than ten $b$-beautiful integers.
+1.
+Understanding the Problem: - We need to identify $b$-beautiful integers, which have exactly two digits in base $b$.
+- These integers $n$ must satisfy the condition that the sum of their digits equals $\sqrt{n}$.
+6.
+Conclusion: - The least integer $b\geq 2$ for which there are more than ten $b$-beautiful integers is $b=13$.
+Therefore, the smallest $b$ such that there are more than ten $b$-beautiful integers is $b=13$.
+Next, let’s write the corresponding formal proof in Lean 4 to prove this.
+# Formal proof in Lean 4:
+‘‘‘lean4
+theorem lean_workbook_plus_1377 (b : ℕ) (hb : 2 ≤ b)
+(h : (∃ n, (Nat.sqrt n = √n ∧ Nat.digits b n = [n / b, n % b]))
+  → ¬ (∃ n, (Nat.sqrt n = √n ∧ Nat.digits b n = [n / b, n % b]))) : 13 ≤ b :=
+  by
+  contrapose! h
+  interval_cases b <;> norm_num [Nat.digits_eq_of_eq_mul_left, Nat.digits_eq_of_eq_mul_right]
+‘‘‘lean_output
+27.000000
+Let’s use Python code to perform these calculations. # Code in Python:
+‘‘‘python
+import sympy as sp
+import math
+def is_b_beautiful(n, b):
+\"\"\"Check if a number n is b-beautiful.\"\"\"
+digits = [n // b, n % b]
+return sum(digits) == math.sqrt(n)
+def count_b_beautiful_numbers(b):
+\"\"\"Count the number of b-beautiful numbers for a given b.\"\"\"
+count = 0
+for n in range(1, b * b):
+if is_b_beautiful(n, b):
+count += 1
+return count
+def find_smallest_b_with_more_than_ten_beautiful_numbers():
+\"\"\"Find the smallest b >= 2 such that there are more than ten b-beautiful numbers.\"\"\"
+b = 2
+while True:
+if count_b_beautiful_numbers(b) > 10:
+return b
+b += 1
+# Find the smallest b
+smallest_b = find_smallest_b_with_more_than_ten_beautiful_numbers()
+print(smallest_b)
+‘‘‘output
+211
+The smallest integer $b\geq 2$ for which there are more than ten $b$-beautiful integers is $b=211$. Thus, the answer is $\boxed{211}$.
+Figure 12: A case of self-correction behaviors. Some lengthy content is omitted with "…" for brevity.